### v4

Abstract
For $p\in[1,\infty]$, we define a smooth manifold structure on the set $AC_{L^{p}}([a,b],N)$ of absolutely continuous functions $\gamma\colon[a,b]\to N$ with $L^{p}$-derivatives for all real numbers $a<b$ and each smooth manifold $N$ modeled on a sequentially complete locally convex topological vector space, such that $N$ admits a local addition. Smoothness of natural mappings between spaces of absolutely continuous functions is discussed, like superposition operators $AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2})$, $\eta\mapsto f\circ\eta$, for a smooth map $f\colon N_{1}\to N_{2}$. Let $G$ be a right half-Lie group. For $1\leq p<\infty$, we show certains properties of $L^{p}$-semiregularity of $G$ under certains condicions on $G$. For $r\in\mathbb{N}$, we prove that the right half-Lie groups $\text{Diff}_{K}^{r}(\mathbb{R})$ and $\text{Diff}^{r}(M)$ are $L^{p}$-semiregular. Here $K$ is a compact subset of $\mathbb{R}$ and $M$ is a compact smooth manifold (possible with boundary). An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\text{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$. For the preceding examples, the evolution map Evol is continuous.
Manifolds of absolutely continuous functions with values in
an infinite-dimensional manifold and regularity
properties of half-Lie groups
Matthieu F. Pinaud
MSC 2020 subject classification: 58D25 (primary); 22E65, 34A12, 46T10, 58D05.
Keywords: Infinite-dimensional manifolds, manifold of mappings, absolutely continuous functions, half-Lie group, regular Lie group, regularity, diffeomorphism group, evolution, initial value problem.
1 Introduction
The study of mapping manifolds is an important branch of geometry and global analysis. For example, for $k\in\mathbb{N}\cup\{0\}$, a well-known case is the manifold of the set of $C^{k}$-maps $f:M\to N$, denoted by $C^{k}(M,N)$, for a compact manifold $M$ and a smooth manifold $N$ admitting local addition (see e.g. [1, 9, 28]). In [24], Krikorian shows how to contruct smooth manifold structures for set of functions with Hölder conditions and Sobolev functions.
Our main focus will be the study of absolutely continuous functions with values in a manifold $N$ modeled on a sequentially complete locally convex space which admits a local addition. For a Hilbert manifold $N$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [10] and [22], cf. also [39]). For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, Pierron and Trouvé show that a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [24], which is similar to Palais’ use of Banach section functors [35]. A Lie group structure (and hence a smooth manifold structure) on $AC_{L^{p}}([0,1],G)$ was obtained in [33] for each Lie group $G$ modeled on a sequentially complete locally convex space (as in [30]), generalizing the case of Fréchet–Lie groups treated in [14].
We construct manifolds of absolutely continuous functions in higher generality. To formulate the main result, let us fix notation.
Definition 1.1
Let $N$ be a smooth manifold modeled on real a locally convex space, with tangent bundle $TN$ and its bundle projection $\pi_{TN}\colon TN\to N$. A local addition on $N$ is a map
$\Sigma\colon\Omega\to N,$
defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $p\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
$\theta\colon\Omega\to N\times N,\quad v\mapsto(\pi_{TN}(v),\Sigma(v))$
has open image and is a $C^{\infty}$-diffeomorphism onto its image.
Throughout the following, $a<b$ are real numbers and $p\in[1,\infty]$. For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.12 and 2.23, respectively (see also [33]). For $\eta\in AC_{L^{p}}([a,b],N)$, the pointwise operations make
$\Gamma_{AC}(\eta):=\{\tau\in AC_{L^{p}}([a,b],TN)\colon\pi_{TN}\circ\tau=\eta\}$
a vector space; we endow it with a natural topology making it a locally convex topological vector space (cf. Definition 2.26). We shall see that
${\mathcal{V}}_{\eta}:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
is an open 00-neighborhood in $\Gamma_{AC}(\eta)$. Setting
${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(\eta(t),\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
the map
$\Psi_{\eta}\colon{\mathcal{V}}_{\theta}\to{\mathcal{U}}_{\theta},\quad\tau\mapsto\Sigma\circ\tau$
is a bijection (see Remark 3.3). We show (see Theorem 3.5):
Theorem 1.2
For each smooth manifold $N$ modeled on a sequentially complete locally convex space and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ of all $AC_{L^{p}}$-maps $\gamma\colon[a,b]\to N$ admits a smooth manifold structure such that for each local addition $\Sigma\colon\Omega\to N$, the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $C^{\infty}$-diffeomorphism.
Using the smooth manifold structures just described, we find:
Theorem 1.3
Let $f\colon N_{1}\to N_{2}$ be a smooth map between smooth manifolds $N_{1}$ and $N_{2}$ modeled on sequentially complete locally convex spaces such that $N_{1}$ and $N_{2}$ admit a local addition and $p\in[1,\infty]$. We then have $f\circ\gamma\in AC_{L^{p}}([a,b],N_{2})$ for all $\gamma\in AC_{L^{p}}([a,b],N_{1})$; the map
$AC_{L^{p}}([a,b],f)\colon AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2}),\quad\gamma\mapsto f\circ\gamma$
is smooth.
More generally, $AC_{L^{p}}([a,b],f)$ is $C^{r}$ for $r\in\mathbb{N}\cup\{0,\infty\}$ whenever $f$ is $C^{r+2}$ (see Proposition 3.8).
If, in the situation of Definition 1.1, $N$ is a $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition. In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{AC}(\eta)$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 3.6). If $N_{j}$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex topological $\mathbb{K}$-vector space such that $N_{j}$ admits a $\mathbb{K}$-analytic local addition for $j\in\{1,2\}$, then the map $AC_{L^{p}}([a,b],f)$ described in Theorem 1.3 is $\mathbb{K}$-analytic for all $p\in[1,\infty]$ (see Corollary 3.9).
If $\pi_{TN}:TN\to N$ denotes the tangent bundle of $N$. Followings results on more general manifolds of mappings (see e.g. [3]), we can conclude that the map
$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
defines a vector bundle on $AC_{L^{p}}([a,b],N)$. The proofs are analogous to the reference given, but we will show them for the sake of completeness.
Manifolds of absolutely continuous paths occur in the regularity theory of infinite-dimensional Lie groups. Consider a Lie group $G$ modeled on a sequentially complete locally convex space, with tangent space ${\mathfrak{g}}:=T_{e}G$ at the neutral element $e\in G$. For $g\in G$, let $\rho_{g}\colon G\to G$, $x\mapsto xg$ be the right multiplication with $g$. Then
$TG\times G\to TG,\quad(v,g)\mapsto T\rho_{g}(v)=:v.g$
is a smooth map and a right action of $G$ on $TG$. The following concept strengthens “regularity” in the sense of Milnor [30].
Definition 1.4
Following [33] (cf. also [14]), $G$ is called $L^{p}$-semiregular if, for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$, the initial value problem
$\dot{\eta}(t)$ $=\gamma(t).\eta(t),\quad t\in[0,1]$ (1.1)
$\eta(0)$ $=e$ (1.2)
has an $AC_{L^{p}}$-solution $\eta\colon[0,1]\to G$. Then $\eta$ is necessarily unique; we call $\eta$ the evolution of $\gamma$ and write $\operatorname{Evol}([\gamma]):=\eta$. If $G$ is $L^{p}$-semiregular and $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$ is smooth, then $G$ is called $L^{p}$-regular.
Remark 1.5
If $G$ is modeled on a Fréchet space, (1.1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure. In the general case, $\eta$ is required to be a Carathéodory solution to (1.1), i.e., it solves the corresponding integral equation piecewise in local charts. We mention that a Lie group $G$ is $L^{p}$-regular if and only if $G$ is $L^{p}$-semiregular and $\operatorname{Evol}$ is smooth as map
$L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$
(see [33]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [15, Theorem E]).
Now consider a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$. Thus $G$ is a group, endowed with a smooth manifold structure modeled on $E$ which makes $G$ a topological group and turns the right translation $\rho_{g}\colon G\to G$ into a smooth mapping for each $g\in G$ (cf. [6, 27]). Let us use notation as in the case of Lie groups.
Definition 1.6
We say that a half-Lie group $G$ is $L^{p}$-semiregular if the differential equation
$\dot{y}(t)=\gamma(t).y(t),\quad t\in[0,1]$ (1.3)
satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [18]) and the initial value problem (1.1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
The Lie group $\text{Diff}(M)$ of $C^{\infty}$-diffeomorphisms of a compact smooth manifold $M$ without boundary is known to be $L^{1}$-regular, and also the Lie group $\text{Diff}_{K}(\mathbb{R}^{n})$ of all $C^{\infty}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\phi(x)=x$ for all $x\in\mathbb{R}^{n}\setminus K$, for each compact subset $K\subseteq\mathbb{R}^{n}$ (see [14]). For each positive integer $r$, the following analogues are obtained (see Theorems  6.5 and 5.3):
Theorem 1.7
Let $1\leq p<\infty$. For each compact smooth manifold $M$ without boundary and $r\in\mathbb{N}$, the half-Lie group $G:=\mbox{\emph{Diff}}^{r}(M)$ of all $C^{r}$-diffeomorphisms $\phi\colon M\to M$ is $L^{p}$-semiregular. Moreover, its evolution map
$\mbox{\emph{Evol}}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$
is continuous.
Here $\mathfrak{g}$ is the Banach space of $C^{r}$-vector fields on $M$.
Theorem 1.8
Let $1\leq p<\infty$ and $r\in\mathbb{N}$. For each positive integer $n$ and compact subset $K$ of $\mathbb{R}^{n}$, the half-Lie group $G:=\mbox{\emph{Diff}}^{r}_{K}(\mathbb{R}^{n})$ of all $C^{r}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ with $\phi|_{\mathbb{R}^{n}\setminus K}=\text{id}_{\mathbb{R}^{n}\setminus K}$ is $L^{p}$-semiregular. Moreover, its evolution map
$\mbox{\emph{Evol}}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the preceding theorem remains valid.
Here $\mathfrak{g}$ is the Banach space of all $C^{r}$-vector fields on $\mathbb{R}^{n}$ which vanish outside $K$.
For an $L^{p}$-semiregular half-Lie group $G$ admitting a local addition with $1\leq p<\infty$, the smooth manifold structure on $AC_{L^{p}}([0,1],G)$ provided by Theorem 1.2 makes it possible to discuss continuity properties and differentiability properties of the evolution map as a mapping
$\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G).$
So far, we have one positive result in this regard:
Theorem 1.9
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space space $E$ which admits a local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G),\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
If the restriction of the right action
$\tau:T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g$
is continuous, then the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
So far, $C^{0}$-regularity has been investigated for the half-Lie group $\text{Diff}^{r}(M)$ in a suitable sense (see the sketch in [29]). Independently, related questions of regularity have been considered by Pierron and Trouvé (see [36]).
2 Preliminaries
Definition 2.1
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be open and $f:U\to F$ be a map. We say that $f$ is $C^{0}$ if it is continuous. We say that $f$ is $C^{1}$ if $f$ is continuous, the directional derivative
$df(x,y):=(D_{y}f)(x):=\lim_{t\to 0}\frac{1}{t}(f(x+ty)-f(x))$
(with $t\neq 0$) exists in $F$ for all $(x,y)\in U\times E$, and $df:U\times E\to F$ is continuous. Recursively, for $k\in\mathbb{N}$ we say that $f$ is $C^{k}$ if $f$ is $C^{1}$ and $df:U\times E\to F$ is $C^{k-1}$. We say that $f$ is $C^{\infty}$ (or smooth) if $f$ is $C^{k}$ for each $k\in\mathbb{N}$.
Definition 2.2
Let $E_{1},E_{2}$ and $F$ be real locally convex spaces, $U\subseteq E_{1}$ and $V\subseteq E_{2}$ be open subsets, $r,s\in\mathbb{N}\cup\{0,\infty\}$ and $f:U\times V\to F$ be a map. If the iterated directional derivatives
$d^{(i,j)}f((x,a),y_{1},...,y_{i},b_{1},...,b_{j}):=(D_{(y_{1},0)}...D_{(y_{i},0)}D_{(0,b_{1})}...D_{(0,b_{j})}f)(x,a)$
exist for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$, and, we assume that the mappings
$d^{(i,j)}f:U\times V\times E_{1}^{i}\times E_{2}^{j}\to F$
are continuous, then $f$ is called a $C^{r,s}$-map.
Definition 2.3
Let $X$ be a locally compact topological space, endowed with a measure
$\mu:\mathcal{B}(X)\to[0,\infty]$ on its $\sigma$-algebra of Borel sets and let $Y$ be a topological space. A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$, there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
For details of the construction of Lebesgue spaces, we refer the reader to see [33].
Definition 2.4
Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue-Borel measure on $\mathbb{R}$. We define the set $\mathcal{L}^{p}([a,b],E)$ as the set of all Lusin $\lambda$-measurable functions $\gamma:[a,b]\to E$ such that for each continuous seminorm $q$ on $E$ we have
$q\circ\gamma\in\mathcal{L}^{p}([a,b],\mathbb{R}).$
And we endow it with the locally convex topology defined by the family of seminorms
$\lVert\cdot\lVert_{\mathcal{L}^{p},q}:\mathcal{L}^{p}([a,b],E)\to[0,\infty[,\quad\lVert\gamma\lVert_{\mathcal{L}^{p},q}:=\lVert q\circ\gamma\lVert_{\mathcal{L}^{p}}.$
Let $\gamma\sim\eta$ if and only if $\gamma(t)=\eta(t)$ for almost all $t\in[a,b]$ and write $[\gamma]$ for the equivalence class of $\gamma$. We define the Hausdorff locally convex space
$L^{p}([a,b],E):=\mathcal{L}^{p}([a,b],E)/[0]$
with seminorms
$\lVert\cdot\lVert_{L^{p},q}:L^{p}([a,b],E)\to[0,\infty[,\quad\lVert[\gamma]\lVert_{L^{p},q}:=\lVert\gamma\lVert_{\mathcal{L}^{p},q}.$
Definition 2.5
Let $E$ be a locally convex space, $a<b$ be real numbers and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue measure on $\mathbb{R}$. We define the set $\mathcal{L}^{\infty}([a,b],E)$ of all Lusin $\lambda$-measurable, essentially bounded functions $\gamma:[a,b]\to E$. For each continuous seminorm $q:E\to\mathbb{R}^{n}$ we define the seminorm
$\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
We endow $\mathcal{L}^{\infty}([a,b],E)$ with the Hausdorff locally convex topology given by these seminorms.
Let $\gamma\sim\eta$ if and only if $\gamma=\eta$ a.e. We define the Hausdorff locally convex space
$L^{\infty}([a,b],E):=\mathcal{L}^{\infty}([a,b],E)/[0]$
with seminorms
$\lVert\cdot\lVert_{L^{\infty},q}:L^{\infty}([a,b],E)\to[0,\infty[,\quad\lVert[\gamma]\lVert_{L^{\infty},q}:=\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}.$
We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ such that the image $\text{Im}(\gamma)$ has compact and metrizable closure, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$. Thus
$L_{rc}^{\infty}([a,b],E):=\mathcal{L}_{rc}^{\infty}([a,b],E)/[0]$
is a Hausdorff locally convex space.
Remark 2.6
Let $E$ be a Frechet space and $p\in[1,\infty]$. If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [14]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([33, Proposition 2.10]).
Remark 2.7
Let $1\leq q\leq p<\infty$, then
$\mathcal{L}^{\infty}([a,b],E)\subseteq\mathcal{L}^{p}([a,b],E)\subseteq\mathcal{L}^{q}([a,b],E)\subseteq\mathcal{L}^{1}([a,b],E)$
as a consequence of Hölder’s inequality.
Definition 2.8
Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$. An element $z\in E$ is called the weak integral of $\gamma$ if
[…]
for each $\forall\alpha\in E^{\prime}$. Then $z$ is called the weak integral of $\gamma$ from $a$ to $b$, and we write $z=:\int_{a}^{b}\gamma(s)ds$.
Definition 2.9
Let $E$ be a locally convex space. We say that a sequence $(x_{n})_{n}\subset E$ is a Cauchy sequence if for each $\varepsilon>0$ and each continuous seminorm $q$ of $E$, there exists an integer $N_{\varepsilon,q}\in\mathbb{N}$, such that for all $m,n\geq N_{\varepsilon,q}$ we have
$q(x_{m}-x_{n})<\varepsilon.$
We say that $E$ is sequentially complete if every Cauchy sequence converge in $E$.
The following lemma [33, Proposition 2.26] allows us to define absolute continuous functions with vector values.
Lemma 2.10
[…]
is continuous.
A related important result of weak integrals is the Mean Value Theorem (see e.g. [13]).
Theorem 2.11
Let $E$ and $F$ be locally convex spaces, $U\subseteq E$ be an open subset, $f:U\to F$ a $C^{1}$-map and $x,y\in U$ such that the line segment $\{tx+(1-t)y\in E:t\in[0,1]\}$ is contained in $U$. Then
[…]
Definition 2.12
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. For $t_{0}\in[a,b]$, we say that a function $\eta:[a,b]\to E$ is $L^{p}$-absolutely continuous (or just absolutely continuous if there is no confusion) if there exists a $[\gamma]\in L^{p}([a,b],E)$ such that
[…]
We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$. Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique (see [32, Lemma 2.28]), the map
$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$ (2.2)
is an isomorphism of vector spaces. We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space (see [32, Definition 3.1]).
The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$.
Lemma 2.13
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. Then the map
$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$ (2.3)
is a linear topological embedding with closed image.
Proof.
We let $I:E\times L^{p}([a,b],E)\to C([a,b],E)$ be the continuous map given by
$I((x,[\gamma])(t):=x+\int_{a}^{t}\gamma(s)ds,\quad t\in[a,b]$
for each $x\in E$ and $[\gamma]\in L^{p}([a,b],E)$. Let $\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E)$, with $t_{0}:=a$, be the isomorphism of topological vector spaces as above. We consider the map
$\Theta:E\times L^{p}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad(x,[\gamma])\mapsto\big{(}I(x,[\gamma]),[\gamma]\big{)}$
which is continuous.
Moreover, since the evaluation map $\varepsilon_{a}:C([a,b],E)\to E$, $\eta\mapsto\eta(a)$ is continuous, the map $\left(\Theta|^{\text{Im}(\Theta)}\right)^{-1}=(\varepsilon_{a},\text{id}_{L^{p}})$ is also continuous. Hence $\Psi$ is a topological embedding.
Let $(\eta_{\alpha},\eta_{\alpha}^{\prime})_{\alpha}$ be a net in $\text{Im}(\Theta)$ that converges to $(\eta,[\gamma])\in C([a,b],E)\times{L^{p}}([a,b],E)$. By continuity of $\varepsilon_{a}$, the net $(\eta_{\alpha}(a))_{\alpha}$ converges to $\eta(a)\in E$, and by continuity of $\Theta$, the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ converges to $\left(I(\eta(a),[\gamma]),[\gamma]\right)\in\text{Im}(\Theta)$. Since the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ also converges to $(\eta,[\gamma])$, we have
$I(\eta(a),[\gamma])=\eta.$
Therefore $\eta^{\prime}=[\gamma]$ and $(\eta,[\gamma])\in\text{Im}(\Theta)$. ∎
Remark 2.14
Let $p\in[1,\infty]$. Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [33, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology. Hence the sets
$AC_{L^{p}}([a,b],V):=\{\eta\in AC_{L^{p}}([a,b],E):\eta([a,b])\subseteq V\}$
are open on $AC_{L^{p}}([a,b],E)$, for each open subset $V\subseteq E$.
For maps between absolute continuous function spaces, we have the following results (see [33]).
Lemma 2.15
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. For $c,d\in\mathbb{R}$ with $c<d$ we define a map $g:[c,d]\to[a,b]$ via
$g(t)=a+\frac{t-c}{d-c}(b-a),\quad t\in[c,d].$
Then $\eta\circ g\in AC_{L^{p}}([c,d],E)$ for each $\eta\in AC_{L^{p}}([a,b],E)$ and the map
$AC_{L^{p}}(g,E):AC_{L^{p}}([a,b],E)\to AC_{L^{p}}([c,d],E),\quad\eta\mapsto\eta\circ g$
is continuous linear.
Lemma 2.16
Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map. Then $f\circ\eta\in AC_{L^{p}}([a,b],F)$ for each $\eta\in AC_{L^{p}}([a,b],V)$.
Lemma 2.17
Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. Let $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{k+2}$-map, then the map
$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
is $C^{k}$. Moreover, we have
$d(f_{*})(\eta,\eta_{1})=df\circ(\eta,\eta_{1})$
for all $(\eta,\eta_{1})\in AC_{L^{p}}([a,b],V)\times AC_{L^{p}}([a,b],E)$.
Remark 2.18
For sequentially complete locally convex spaces $E$ and $F$ we have
$AC_{L^{p}}([a,b],E\times F)\cong AC_{L^{p}}([a,b],E)\times AC_{L^{p}}([a,b],F).$
Definition 2.19
Let $E$ and $F$ be complex topological vector spaces, where is $F$ is locally convex, $U\subseteq E$ be an open subset and $f:U\to F$ be a mapping. We say that $f$ is complex analytic if it is continuous and, for each $x\in U$, there exists a 00-neighbouhood $V\subseteq E$ such that $x+V\subseteq U$ and certain continuous homogeneous polynomials $\beta_{n}:E\to F$ of degree $n$, such that $f$ admits the expansion: $f(x+y)=\sum_{n=0}^{\infty}\beta_{n}(y)$, for all $y\in V$.
For our context, we present an application of [4, Proposition 7.7] to our particular case.
Lemma 2.20
Let $E$ and $F$ be complex locally convex spaces, and $f:U\to F$ be a mapping defined on an open subset of $E$. Then $f$ is complex analytic if and only if $f$ is smooth and the mapping $df(x,\cdot):E\to F$ is complex linear for each $x\in U$.
Definition 2.21
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be an open subset and $f:U\to F$ be a map. We say that $f$ is real analytic if it extends to an analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$, where $E_{\mathbb{C}}$ and $F_{\mathbb{C}}$ denotes the complexification of $E$ and $F$, respectively.
Lemma 2.22
Let $E$ and $F$ be sequentially complete locally convex spaces over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, $p\in[1,\infty]$, $V\subseteq E$ be an open set and $f:V\to F$ be a $\mathbb{K}$-analytic map. Then the map
$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
is $\mathbb{K}$-analytic.
Proof.
First we consider the case $\mathbb{K}=\mathbb{C}$. By Lemma 2.17 the map $f_{*}$ is smooth and $d(f_{*})$ is complex linear in the second variable, hence by Lemma 2.20 the map $f_{*}$ is complex analytic.
If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})_{*}$ is the complex analytic extension of $f_{*}$, whence $f_{*}$ is real analytic. ∎
Definition 2.23
Let $N$ be a smooth manifold modeled with on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, there exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verifies
  • i)
    $\eta([t_{j-1},t_{j}])\subseteq U_{j}$.
  • ii)
    $\varphi_{j}\circ\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],E)$.
In this case, we say that these charts verify the definition of $L^{p}$-absolute continuity for $\eta$. If there is no confusion, we simple call $\eta$ absolutely continuous.
For the case of absolutely continuous functions with values in a manifold, the following facts are available (see [33]).
Lemma 2.24
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $\eta\in AC_{L^{p}}([a,b],N)$, then
$\varphi\circ\eta|_{[\alpha,\beta]}\in AC_{L^{p}}([\alpha,\beta],E),$
for each chart $\varphi:U\to V$ of $N$ and each subinterval $[\alpha,\beta]\subseteq[a,b]$ such that $\eta([\alpha,\beta])\subseteq U$.
Lemma 2.25
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces and $p\in[1,\infty]$. If $f:M\to N$ is a $C^{1}$-map, then $f\circ\eta\in AC_{L^{p}}([a,b],N)$ for each $\eta\in AC_{L^{p}}([a,b],M)$.
Definition 2.26
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its tangent bundle and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$ (2.4)
and we endow it with the pointwise operations, making it a vector space.
Consider a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$. Since $\eta([t_{i-1},t_{i}])\subseteq U_{i}$, we have
$\sigma([t_{i-1},t_{i}])\subseteq TU_{i},\quad\text{for all }\sigma\in{\Gamma_{AC}(\eta)}$
and for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is the initial topology with respect to the linear mappings
$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$ (2.5)
with $i\in\{1,...,n\}$.
Proposition 2.27
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$, then the map
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$ (2.6)
is a linear topological embedding with closed image given by the set of all elements $(\tau_{i})_{i=1}^{n}$ such that
$\tau_{i}(t_{i})=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)},\quad\text{ for all }i\in\{1,...,n-1\}.$
Proof.
The linear map $\Phi_{\eta,P}$ is continuous by the previous definition and it is injective.
Let $i\in\{1,...,n\}$. If $W_{i}=\varphi_{i+1}(U_{i}\cap U_{i+1})$, then the map
$g_{i}:W_{i}\times E\to E,\quad(x,y)\mapsto d\varphi_{i}\circ(T\varphi_{i+1})^{-1}(x,y),$
is continuous and linear in the second component. This enables us to define the closed vector subspace $K$ given by the elements
$(\tau_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)$
which verify
$\tau_{i}(t_{i})=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)},\quad\text{ for all }i\in\{1,...,n-1\}$
We will show that the image of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincides with the closed subspace $K$. Indeed, the space $\text{Im}(\Phi_{\eta,P})$ is contained in $K$ by definition of $\Phi_{\eta,P}$. Let us consider now $\tau=(\tau_{i})_{i=1}^{n}\in K$. We define the maps
$\sigma_{i}:[t_{i-1},t_{i}]\to TN,\quad s\mapsto(T\varphi_{i})^{-1}\Big{(}\varphi_{i}\circ\eta(s),\tau_{i}(s)\Big{)}$
and
$\sigma_{\tau}:[a,b]\to TN,\quad t\mapsto\sigma_{i}(t),\text{ for }t\in[t_{i-1},t_{i}].$
By the Glueing Lemma, the map $\sigma_{\tau}$ is continuous. Moreover, by Lemma 2.25 each function $\sigma_{i}$ is absolutely continuous, hence $\sigma$ it is too. Since ${\pi_{TN}\circ\sigma_{\tau}=\eta}$ we have that $\sigma_{\tau}\in{\Gamma_{AC}(\eta)}$ and $\Phi_{\eta,P}(\sigma_{\tau})=(\tau_{i})_{i=1}^{n}$, and $\text{Im}(\Phi_{\eta,P})=K$.
It remains to show that the inverse map
$\Phi_{\eta,P}^{-1}:\text{Im}(\Phi_{\eta,P})\to\Gamma_{AC}(\eta),\quad(\tau_{i})_{i=1}^{n}\mapsto\sigma_{\tau}$
is continuous. If $h_{i}$ are the functions that define the topology (Definition 2.26), then for each $i\in\{1,...,n\}$ we have $h_{i}\circ\Phi_{\eta,P}^{-1}=q_{i}$, where $q_{i}$ is the continuous linear map
$q_{i}:\prod_{j=1}^{n}AC_{L^{p}}([t_{j-1},t_{j}],E)\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad(\eta_{j})_{j=1}^{n}\mapsto\eta_{i}.$
Hence $\Phi_{\eta,P}^{-1}$ is continuous. ∎
Remark 2.28
From now we consider the map $\Phi_{\eta,P}$ as the homeomorphism
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\text{Im}(\Phi_{\eta,P}).$
Corollary 2.29
Let $N$ be a smooth manifold modeled on a $E$ Banach space (resp. Frechet space), $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then the vector space ${\Gamma_{AC}(\eta)}$ is a Banach space (resp. Frechet space).
Proof.
This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. Frechet space) ∎
Proposition 2.30
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces, $k\in\mathbb{N}\cup\{0,\infty\}$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],M)$. If $f:M\to N$ is a $C^{k+3}$-map, then $Tf\circ\sigma\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Moreover, the map
$\widetilde{f}:\Gamma_{AC}(\eta)\to\Gamma_{AC}(f\circ\eta),\quad\sigma\mapsto Tf\circ\sigma$
is continuous linear.
Proof.
Let $E_{M}$ and $E_{N}$ be the modeling spaces of $M$ and $N$ respectively. By Lemma 2.16 we have $f\circ\eta\in AC_{L^{p}}([a,b],N)$ and $Tf\circ\sigma\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Since $T_{\eta(t)}f\left(\sigma(t)\right)\in T_{f\circ\eta(t)}N$ for each $t\in[a,b]$, we have
$\pi_{TN}\circ(Tf\circ\sigma)=f\circ\eta.$
Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is clear.
Without loss of generality, we can choose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there exist families of charts $\{(\varphi_{i},U_{i}):i\in\{1,...,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, such that
$f(U_{i})\subseteq V_{i},\quad\text{for each }i\in\{1,...,n\}.$
For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\eta_{i}=\eta|_{[t_{i-1},t_{i}]}$ and $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$.
Moreover, for each $i\in\{1,...,n\}$ we define the maps
$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ T\varphi_{i}^{-1}\right)\circ(\varphi_{i}\circ\eta_{i},\tau)$
and
$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\left(\tau_{i}\right)_{i=1}^{n}\mapsto\Big{(}F_{i}(\tau_{i})\Big{)}_{i=1}^{n}$
which are continuous by Lemma 2.17. We will show that
$F\left(\text{Im}\left(\Phi_{\eta,P}\right)\right)\subseteq\text{Im}\left(\Phi_{f\circ\eta,P}\right).$
Let $i\in\{1,..,n-1\}$. We denote
$\tau_{i}:$ $=F_{i}(d\varphi_{i}\circ\sigma_{i})$
$=(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1})(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ Tf\circ\sigma_{i}$
then
$d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\phi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
$=d\phi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
$=\tau_{i}(t_{i})$
Hence $F\circ\Phi_{\eta,P}(\sigma)\in\text{Im}(\Phi_{f\circ\eta,P})$ and in consequence
$\widetilde{f}=\Phi_{f\circ\eta,P}^{-1}\circ F\circ\Phi_{\eta,P}.$
Thus $\widetilde{f}$ is continuous. ∎
Remark 2.31
The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by the previous proposition the map
$\widetilde{\text{id}_{M}}:{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\text{id}_{M}\circ\eta),\quad\sigma\mapsto T\text{id}_{M}\circ\sigma$
is smooth regardless of the partition or charts chosen.
Remark 2.32
For $\eta\in C([a,b],N)$ we endow the vector space
$\Gamma_{C}(\eta)=\{\sigma\in C([a,b],TN):\pi_{TN}\circ\sigma=\gamma\}$
with the compact-open topology. Since each inclusion
$AC_{L^{p}}([t_{i-1},t_{i}],E)\to C([t_{i-1},t_{i}],E)$
is continuous [33, Lemma 3.2], by Proposition 2.27 the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous. This implies that the set
$\mathcal{V}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq V\}$
is open in ${\Gamma_{AC}(\eta)}$ for each open subset $V\subseteq TN$.
Proposition 2.33
Let $N_{1}$ and $N_{2}$ be smooth manifolds modeled on sequentially complete locally convex spaces, $p\in[1,\infty]$ and $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the $i$-projection for $i\in\{1,2\}$. If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$, then the map
$\mathcal{P}:\Gamma_{AC}(\eta_{1},\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
is a linear homeomorphism.
Proof.
By Proposition 2.30 the map $\mathcal{P}$ is continuous and clearly linear. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$; let $\{(\phi_{1,i},U_{1,i})\colon i\in\{1,\ldots,n\}\}$ and $\{(\phi_{2,i},U_{2,i})\colon i\in\{1,\ldots,n\}\}$ be families of charts of $N_{1}$ and $N_{2}$, respectively, that verify the definition of absolute continuity for $\eta_{1}$ and $\eta_{2}$, respectively. Then $\eta:=(\eta_{1},\eta_{2})\colon[a,b]\to N_{1}\times N_{2}$ is $L^{p}$-absolutely continuous and it is clear that the charts $\{(\phi_{1,i}\times\phi_{2,i},U_{1,i}\times U_{2,i})\colon i\in\{1,\ldots,n\}\}$ satisfy the condition of absolute continuity for $\eta$. For $j\in\{1,2\}$, consider the linear topological embedding
$\Phi_{\eta_{j},P}\colon\Gamma_{\eta_{j}}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{j}),\;\;\tau\mapsto(d\phi_{j,i}\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
where $E_{j}$ is the modeling space of $N_{j}$. Also
$\Phi_{\eta,P}\colon\Gamma_{\eta}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),\;\;\tau\mapsto((d\phi_{1,i}\times d\phi_{2,i})\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
is a linear topological embedding; here $T_{(x_{1},x_{2})}(N_{1}\times N_{2})$ is identified with $T_{x_{1}}N_{1}\times T_{x_{2}}N_{2}$. For $i\in\{1,\ldots,n\}$, let
$\alpha_{i}\colon AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\times AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2})$
be the map taking a pair $(f_{1},f_{2})$ of functions to the function $t\mapsto(f_{1}(t),f_{2}(t))$; we know that $\alpha_{i}$ is an isomorphism of topological vector spaces. Then also
$\alpha\colon\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\right)\times\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\right)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),$
$((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})\mapsto(\alpha(f_{i},g_{i}))_{i=1}^{n}$
is an isomorphism of topological vector spaces. If $(f_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{1},P}$ and $(g_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{2},P}$, then $\alpha((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})$ is in the image of $\Phi_{\eta,P}$, as the compatibility at the endpoints can be checked by considering the components in $E_{1}$ and $E_{2}$. We can therefore define a function
$\Theta:=\Phi_{\eta,P}^{-1}\circ\alpha\circ(\Phi_{\eta_{1},P}\times\Phi_{\eta_{2},P})\colon\Gamma_{\eta_{1}}\times\Gamma_{\eta_{2}}\to\Gamma_{\eta},$
which is continuous and linear. We readily check that $\mathcal{P}(\Theta(\sigma,\tau))=(\sigma,\tau)$ for all $\sigma\in\Gamma_{\eta_{1}}$ and $\tau\in\Gamma_{\eta_{2}}$. Hence $\mathcal{P}$ is surjective and thus bijective, with $\mathcal{P}^{-1}=\Theta$ a continuous map. ∎
Proposition 2.34
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. For a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$, the map
$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}\left(\eta|_{[t_{i-1},t_{i}]}\right),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is a linear topological embedding with closed image.
Proof.
Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{j-1},t_{j}\}$, since $\eta_{j}:=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
$\rho=\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}^{-1}\right)\circ\Phi_{\eta,P}.$
The image is given by the closed subspace
$\text{Im}(\rho)=\left\{(\tau_{i})_{i=1}^{n}\in\prod_{i=1}^{n}\Gamma_{AC}(\eta|_{[t_{i-1},t_{i}]}):\tau_{i}(t_{i})=\tau_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\right\}.$
Thus $(\rho|^{\text{Im}(\rho)})^{-1}:\text{Im}(\rho)\mapsto{\Gamma_{AC}(\eta)}$ is well defined and
$\left(\rho|^{\text{Im}(\rho)}\right)^{-1}=\Phi_{\eta,P}^{-1}\circ\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}\right).$
Proposition 2.35
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $g:[c,d]\to[a,b]$ is the map as in Lemma 2.15, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then the map
$L_{g}:{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\eta\circ g),\quad\sigma\mapsto\sigma\circ g$
is continuous linear.
Proof.
Let $\eta\in AC_{L^{p}}([a,b],N)$. We will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $N$ that verify the definition of absolute continuity for $\eta$. Since $g$ is a strictly increasing function, we can define a partition $Q=\{s_{1},...,s_{n}\}$ of $[c,d]$ such that $g(s_{i})=t_{i}$, for each $i\in\{1,...,n\}$. Moreover, given that
$\varphi_{i}\circ\left(\eta\circ g\right)|_{[s_{i-1},s_{i}]}=\left(\varphi_{i}\circ\eta|_{[t_{i-1},t_{i}]}\right)\circ g|_{[s_{i-1},s_{i}]}$
we have $\eta\circ g\in AC_{L^{p}}([a,b],E)$ by Lemma 2.15. Analogously, we have that $\sigma\circ g\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$ and
$\pi_{TN}\circ(\sigma\circ g)=\eta\circ g.$
Hence $L_{g}(\sigma)\in\Gamma_{AC}(\eta\circ g)$. To see the continuity of $L_{g}$, for each $i\in\{1,...,n\}$ we define the maps
$G_{i}:AC_{L^{p}}([t_{i+1},t_{i}],E)\to AC_{L^{p}}([s_{i+1},s_{i}],E),\quad\tau\mapsto\tau\circ g|_{[s_{i+1},s_{i}]}$
which are continuous by Lemma 2.15. Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as in Proposition 2.27) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
$\tau_{i}\circ g(s_{i})$ $=\tau_{i}(t_{i})$
$=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta\circ g(s_{i}),\tau_{i+1}\circ g(s_{i})\Big{)}$
for each $i\in\{1,...,n-1\}$. If $G=(G_{1}\times...\times G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
$L_{g}=\Phi_{\eta\circ g,Q}^{-1}\circ G\circ\Phi_{\eta,Q}.$
Hence $L_{g}$ is continuous and clearly linear. ∎
Proposition 2.36
Let $N$ be a smooth manifold modeled on sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then the evaluation map
$\epsilon:{\Gamma_{AC}(\eta)}\times[a,b]\to TN,\quad(\sigma,t)\mapsto\sigma(t)$
is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$, the point evaluation map
$\epsilon_{t}:{\Gamma_{AC}(\eta)}\to TN,\quad\sigma\mapsto\sigma(t)$
is smooth.
Proof.
The evaluation map
$\tilde{\epsilon}:\Gamma_{C}(\eta)\times[a,b]\to TN,\quad(\sigma,t)\mapsto\sigma(t)$
is continuous and the evaluation map $\tilde{\epsilon}_{t}:\Gamma_{C}(\eta)\to TN$, $\sigma\mapsto\sigma(t)$ is smooth for each $t\in[a,b]$ (see [3]). Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma}\times\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 2.32. ∎
3 Manifolds of absolutely continuous functions
Definition 3.1
Let $N$ be a smooth manifold and $\pi_{TN}:TN\to N$ its tangent bundle. A local addition is a smooth map ${\Sigma:\Omega\to N}$ defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section ${0_{N}:=\{0_{p}\in T_{p}N:p\in N\}}$ such that
  • a)
    $\Sigma(0_{p})=p$ for all $p\in N$.
  • b)
    The image $\Omega^{\prime}:=\big{(}\pi_{TN},\Sigma\big{)}(\Omega)$ is open in $N\times N$ and the map
    $\theta_{N}:\Omega\to\Omega^{\prime},\quad v\mapsto\big{(}\pi_{TN}(v),\Sigma(v)\big{)}$
    is a $C^{\infty}$-diffeomorphism.
Moreover, if $T_{0_{p}}(\Sigma|_{T_{p}N})=id_{T_{p}N}$ for all $p\in N$, we say that the local addition $\Sigma$ is normalized. We denote the local addition as the pair $(\Omega,\Sigma)$.
If $\theta_{N}:\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, we call ${\Sigma:\Omega\to N}$ a $\mathbb{K}$-analytic local addition.
Remark 3.2
Every paracompact finite-dimensional smooth manifold admits a local addition. If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [3, Lemma A.11]. Moreover, each manifold which admits a local addition also admits a normalized local addition [3, Lemma A.14]. From now we will assume that each local addition is normalized.
Remark 3.3
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. We define the sets
$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}$
which is open in ${\Gamma_{AC}(\eta)}$ by Remark 2.32 and
$\mathcal{U}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N):(\eta,\gamma)([a,b])\subseteq\Omega^{\prime}\}.$
Lemma 2.25 enable us to define the map
$\Psi_{\eta}:=AC_{L^{p}}([a,b],\Sigma):\mathcal{V}_{\eta}\to\mathcal{U}_{\eta},\quad\sigma\mapsto\Sigma\circ\sigma.$
with inverse given by
$\Psi_{\eta}^{-1}:\mathcal{U}_{\eta}\to\mathcal{V}_{\eta},\quad\gamma\mapsto\theta_{N}^{-1}\circ(\eta,\gamma).$
The following lemma is just an application of [4, Lemma 10.1] to our particular case.
Lemma 3.4
Let $E$ and $F$ sequentially complete locally convex spaces, $U\subseteq E$ open and $f:U\to F$ a map. If $F_{0}\subseteq F$ is a closed vector subspace and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
Theorem 3.5
For each smooth manifold $N$ modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ admits a smooth manifold structure such that the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $C^{\infty}$-diffeomorphism.
Proof.
We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps
$\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta},\quad\text{for each }\eta\in AC_{L^{p}}([a,b],N).$
If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as in Remark 3.3, then the final topology on $C([a,b],N)$ with respect to this maps coincides with its compact-open topology. Hence the inclusion map
$J:AC_{L^{p}}([a,b],N)\to C([a,b],N),\quad\gamma\mapsto\gamma$
is continuous. Moreover, for each $\eta\in AC_{L^{p}}([a,b],N)$ the set
$\mathcal{U}_{\eta}^{C}:=\{\gamma\in C([a,b],N):(\eta,\gamma)([a,b])\subseteq\Omega^{\prime}\}$
is open in $C([a,b],N)$, whence
$\mathcal{U}_{\eta}=\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],N)$
is open in $AC_{L^{p}}([a,b],N)$.
The goal is to make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$. We need to show that the charts are compatible, i.e., the smoothness of the map
$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma)$
for each $\eta,\xi\in AC_{L^{p}}([a,b],N)$ such that the open set
$\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})=\left(\Psi_{\eta}^{C}\right)^{-1}(\mathcal{U}_{\eta}^{C}\cap\mathcal{U}_{\xi}^{C})\cap{\Gamma_{AC}(\eta)}$
is not empty.
Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...,n\}\}$ be charts that verify the definition of absolute continuity for $\eta$ and $\xi$, respectively. For each $\sigma\in AC_{L^{p}}([a,b],TN)$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$ and
$\Lambda_{\xi,\eta}(\sigma_{i}):=\Lambda_{\xi,\eta}(\sigma)\big{|}_{[t_{i-1},t_{i}]}.$
We will study the smoothness of the composition
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$ $\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)$
$\sigma$ $\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
that by Lemma 3.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$, we denote $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and $\xi_{i}:=\xi|_{[t_{i-1},t_{i}]}$ and we have
$d{\phi}_{i}\circ\Lambda_{\xi,\eta}|_{[t_{i-1},t_{i}]}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi_{i},\Sigma\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)).$
Because all of the functions involved are continuous and have an open domain, also the composition
$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z))$ (3.1)
has an open domain $\mathcal{O}_{i}$. Hence the map $H_{i}:\mathcal{O}_{i}\to E$ is smooth. By Lemma 2.17, the map
$AC_{L^{p}}([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
is smooth. Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.18), if we fix the functions $\phi_{i}\circ\xi_{i}$ and $\varphi_{i}\circ\eta_{i}$, we have the continuous linear map
$AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E),\quad\tau\mapsto(\phi_{i}\circ\xi_{i},\varphi_{i}\circ\eta_{i},\tau).$
We write $W_{i}$ for the preimage of $AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})$ under this map. Then the map
$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
is also smooth. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\sigma\mapsto d\phi_{i}\circ\sigma_{i}$ are continuous by definition of the topology, rewriting we have
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}(\sigma)=\left(\Theta_{i}\circ h_{i}(\sigma)\right)_{i=1}^{n}$
for each $\sigma\in\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$, hence $\Lambda_{\xi,\eta}$ is smooth. ∎
Proceding in the same way, using the fact that compositions of $\mathbb{K}$-analytic maps are $\mathbb{K}$-analytic and using the analytic version of Lemma 3.4 (see [13]), we obtain the analogous case.
Corollary 3.6
For each $\mathbb{K}$-analytic manifold $N$ modeled on a sequentially complete locally convex space $E$ which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ admits a $\mathbb{K}$-analytic manifold structure such that the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $\mathbb{K}$-analytic diffeomorphism.
Proposition 3.7
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space which admits a local addition and $1\leq q\leq p\leq\infty$. Then
$AC_{L^{\infty}}([a,b],N)\subseteq AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)\subseteq AC_{L^{1}}([a,b],N)$
with smooth inclusion maps.
Proof.
Let $\eta\in AC_{L^{p}}([a,b],N)$. Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$.
By [33, Remark 3.2], we know that each inclusion map $AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{q}}([t_{i-1},t_{i}],E)$ is continuous linear, hence $AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)$. By Proposition 2.27, the inclusion is smooth since each map
$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],E),\quad(\eta_{i})_{i=1}^{n}\mapsto(\eta_{i})_{i=1}^{n}.$
is continuous linear. ∎
Proposition 3.8
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. If $f:M\to N$ is a $C^{k+2}$-map, then the map
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is $C^{k}$.
Proof.
Let $E_{M}$ and $E_{N}$ be the modeleding space of $M$ and $N$, respectively. Let $(\Omega_{M},\Sigma_{M})$ and $(\Omega_{N},\Sigma_{N})$ be local additions on $M$ and $N$ respectively. The map makes sense by Lemma 2.25. Let $\eta\in AC_{L^{p}}([a,b],M)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],M)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$, respectively. We see that the set
$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)={\Gamma_{AC}(\eta)}\cap\left(\Psi_{\eta}^{C}\right)^{-1}\left(\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta}^{C})\right)$
is open in ${\Gamma_{AC}(\eta)}$. If $\theta_{N}=(\pi_{TN},\Sigma_{N})$, then we define
$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$.
Proceeding as the proof of Theorem 3.5, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$, $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and
$F(\sigma)|_{[t_{i-1},t_{i}]}:=F(\sigma_{i}).$
We will study the continuity of the map
$\Phi_{f\circ\eta}\circ F:\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)\to\text{Im}(\Phi_{f\circ\eta}),\quad\sigma\mapsto\left(d\phi_{i}\circ F(\sigma_{i})\right)_{i=1}^{n}$
where $\Phi_{f\circ\eta}$ is the topological embedding as in Proposition 2.27. For each $i\in\{1,...,n\}$ and $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$ we have
$d\phi_{i}\circ F(\sigma)|_{[t_{i-1},t_{i}]}$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta_{i},f\circ\Sigma_{M}\circ\sigma_{i}\Big{)}$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta_{i},f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)\Big{)}.$
Because all of the functions involved are continuous and have an open domain, also the composition
$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$ (3.2)
has an open domain $\mathcal{O}_{i}$ in $E_{N}\times E_{M}\times E_{M}$. Hence the map $H_{i}:\mathcal{O}_{i}\to E_{N}$ is smooth. Fixing the absolutely continuous functions $\phi_{i}\circ f\circ\eta_{i}$ and $\varphi_{i}\circ\eta_{i}$, we consider the continuous linear map
$AC_{L^{p}}([t_{i-1},i],E_{M})\to AC_{L^{p}}([t_{i-1},i],E_{N}\times E_{M}\times E_{M}),\quad\tau\mapsto(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
and we write $W_{i}$ for the preimage of $AC_{L^{p}}([a,b],\mathcal{O}_{i})$ under this map. Then the map
$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta_{i},\varphi_{i}\circ\eta_{i},\tau)$
is $C^{k}$. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $\sigma\mapsto d\varphi_{i}\circ\sigma_{i}$ are continuous linear by definition, rewriting we have
$\Phi_{f\circ\eta}\circ F(\sigma)=\left(\Theta_{i}\circ h_{i}(\sigma)\right)_{i=1}^{n}$
for each $\sigma\in\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$, hence the map $AC_{L^{p}}([a,b],f)$ is $C^{k}$. ∎
Proceeding in the same way we have the analogous case.
Corollary 3.9
Let $f:M\to N$ be a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admit $\mathbb{K}$-analytic local additions and $p\in[1,\infty]$. Then the map
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is $\mathbb{K}$-analytic.
Remark 3.10
The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions coincide. Indeed, since the identity map $\text{id}_{N}:N\to N$ is smooth regardless of the chosen local addition, the map
$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{N}\circ\eta$
is, again, smooth regardless of the chosen local addition in each space.
Remark 3.11
The inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth. Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta}^{C},\left(\Psi_{\eta}^{C}\right)^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\in C([a,b],N)$ respectively, then
$\Psi_{\eta}^{-1}\circ J\circ\Psi_{\eta}^{C}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{\eta}^{C})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
is a restriction of the inclusion map ${\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$.
Moreover, if $U\subseteq N$ be a open subset, then the manifold structure induced by $AC_{L^{p}}([a,b],N)$ on the open subset
$AC_{L^{p}}([a,b],U):=\{\eta\in AC_{L^{p}}([a,b],N):\eta([a,b])\subseteq U\}.$
coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$.
Proposition 3.12
Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex spaces which admit a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$. Then the map
$\mathcal{P}:AC_{L^{p}}([a,b],N_{1}\times N_{2})\to AC_{L^{p}}([a,b],N_{1})\times AC_{L^{p}}([a,b],N_{2}),\quad\eta\mapsto(\operatorname{pr}_{1},\operatorname{pr}_{2})\circ\eta$
is a diffeomorphism.
Proof.
By the previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local addition on $N_{1}\times N_{2}$ is
$\Sigma:=\Sigma_{1}\times\Sigma_{2}:\Omega_{1}\times\Omega_{2}\to N_{1}\times N_{2}$
where $\Omega_{1}\times\Omega_{2}\subseteq TN_{1}\times TN_{2}\cong T(N_{1}\times N_{2})$. The map $\mathcal{P}$ is smooth as consequence of the smoothness of the maps
$AC_{L^{p}}([a,b],\operatorname{pr}_{i}):AC_{L^{p}}([a,b],N_{1}\times N_{2})\to AC_{L^{p}}([a,b],N_{i}),$
for each ${i\in\{1,2\}}$.
Let $(\mathcal{U}_{\eta_{1}}\times\mathcal{U}_{\eta_{2}},\Psi_{\eta_{1}}^{-1}\times\Psi_{\eta_{2}}^{-1})$ and $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ be charts in $(\eta_{1},\eta_{2})\in AC_{L^{p}}([a,b],N_{1})\times AC_{L^{p}}([a,b],N_{2})$ and $\mathcal{P}^{-1}(\eta_{1},\eta_{2})=\eta\in AC_{L^{p}}([a,b],N_{1}\times N_{2})$ respectively. Since the map
$\mathcal{Q}:\Gamma_{AC}(\eta)\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{2}),\quad\tau\mapsto(\text{q}_{1},\text{q}_{2})\circ\tau$
is an isomorphism of topological vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
$\Psi_{\eta}^{-1}\circ\mathcal{P}^{-1}\circ(\Psi_{\eta_{1}}\times\Psi_{\eta_{2}})(\sigma_{1},\sigma_{2})$ $=(\pi_{N_{1}\times N_{2}},\Sigma)^{-1}\circ\left(\eta,\mathcal{P}^{-1}\circ(\Sigma_{1}\times\Sigma_{2})(\sigma_{1},\sigma_{2})\right)$
$=(\pi_{N_{1}\times N_{2}},\Sigma)^{-1}\circ\left(\eta,\Sigma\circ\mathcal{Q}^{-1}(\sigma_{1},\sigma_{2})\right)$
$=\mathcal{Q}^{-1}(\sigma_{1},\sigma_{2})$
for all $(\sigma_{1},\sigma_{2})\in(\Psi_{\eta_{1}}^{-1}\times\Psi_{\eta_{2}}^{-1})\left(\mathcal{U}_{\eta_{1}}\times\mathcal{U}_{\eta_{2}}\cap\mathcal{P}(\mathcal{U}_{\eta})\right)$. Hence $\mathcal{P}^{-1}$ is smooth. ∎
Proposition 3.13
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$. For a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ the map
$T:AC_{L^{p}}([a,b],N)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N),\quad\eta\mapsto\left(\eta|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is smooth and a smooth diffeomorphism onto a submanifold of $\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$.
Proof.
It is clear that the map $T$ is well defined and injective. Let $\text{Im}(T)$ be the image of the map $T$. Then
$\text{Im}(T)=\{(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N):\gamma_{i}(t_{i})=\gamma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
Let $\gamma:=(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$. For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts around $\gamma_{i}$. Then the map
$\Psi_{\gamma}^{-1}:=\prod_{i=1}^{n}\Psi_{\gamma_{i}}^{-1}:\prod_{i=1}^{n}\mathcal{U}_{i}\to\prod_{i=1}^{n}\mathcal{V}_{i}$
is a chart around $\gamma$. Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}^{-1}\circ T\circ\Psi_{\eta}$ is just restriction of the product of the restrictions of the smooth maps
${\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta_{i}),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
thus $T$ is a smooth. Now we will show that the image $\text{Im}(T)$ is a submanifold.
Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and
$\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
$\Psi_{\gamma_{i}}^{-1}(\xi_{i})(t_{i})$ $=\theta_{N}^{-1}\circ(\gamma_{i},\xi_{i})(t_{i})$
$=\theta_{N}^{-1}\circ(\gamma_{i}(t_{i}),\xi_{i}(t_{i}))$
$=\theta_{N}^{-1}\circ(\gamma_{i+1}(t_{i}),\xi_{i+1}(t_{i}))$
$=\Psi_{\gamma_{i+1}}^{-1}(\xi_{i+1})(t_{i}).$
This implies that if $K$ denotes the vector space
$K:=\{\left(\sigma_{i}\right)_{i=1}^{n}\in\prod_{i=1}^{n}\Gamma_{AC}(\gamma_{i}):\sigma_{i}(t_{i})=\sigma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold and the map $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$ a diffeomorphism. ∎
Remark 3.14
Let $N$ be a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Since every continuous linear operator is analytic, the isomorphism
${\Gamma_{AC}(\eta)}\to K,\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is $\mathbb{K}$-analytic, which implies that $T$ in Proposition 3.13 is a $\mathbb{K}$-analytic diffeomorphism onto the submanifold $\text{Im}(T)$.
Proposition 3.15
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If $g:[c,d]\to[a,b]$ is a map as in Lemma 2.15, then the map
$AC_{L^{p}}(g,N):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([c,d],N),\quad\eta\mapsto\eta\circ g$
is smooth.
Proof.
By Proposition 2.35 we know that the map is well defined. Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
$\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ(\eta\circ g,\Sigma\circ(\sigma\circ g))$
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$. This set coincides with
$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap C(g,N)^{-1}(\mathcal{U}_{\eta\circ g}^{C})\right)$
which is open given by the continuity of the map $C(g,N)$. Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$. Then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, now $\tau\in\Gamma_{AC}(\alpha)$ and
$\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)$ $=\theta_{N}^{-1}\circ(\alpha,\Sigma\circ\tau)$
$=\Psi_{\alpha}^{-1}\circ\Psi_{\alpha}(\tau)$
$=\tau$
$=\sigma\circ g.$
Hence, $\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}$ is a restriction of the map
${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\mapsto\sigma\circ g$
which is continuous linear by Proposition 2.35. ∎
Proposition 3.16
Let $M$, $N$ and $L$ be smooth manifolds modeled on sequentially complete locally convex spaces which admit a local addition and $p\in[1,\infty]$. If $f:L\times M\to N$ is a $C^{k+2}$-map and $\gamma\in AC_{L^{p}}([a,b],L)$ is fixed, then
$f_{*}:AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ(\gamma,\eta)$
is a $C^{k}$-map.
Proof.
Define the smooth map
$C_{\gamma}:AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],L)\times AC_{L^{p}}([a,b],N),\quad\eta\mapsto(\gamma,\eta)$
Identifying $AC_{L^{p}}([a,b],L)\times AC_{L^{p}}([a,b],N)$ with $AC_{L^{p}}([a,b],L\times N)$, we have
$f_{*}=AC_{L^{p}}([a,b],f)\circ C_{\gamma}.$
Hence $f_{*}$ is $C^{k}$. ∎
Proposition 3.17
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $t\in[a,b]$. Then the evaluation map
$\varepsilon:AC_{L^{p}}([a,b],N)\times[a,b]\to N,\quad(\eta,t)\mapsto\eta(t)$
is continuous and the point evaluation map
$\varepsilon_{t}:AC_{L^{p}}([a,b],N)\to N,\quad\eta\mapsto\eta(t)$
is smooth.
Proof.
The evaluation map
$\varepsilon_{c}:C([a,b],N)\times[a,b]\to N,\quad(\eta,t)\mapsto\eta(t)$
is continuous and the point evaluation $(\varepsilon_{c})_{t}:C([a,b],N)\to N$, $\eta\mapsto\eta(t)$ is smooth for each $t\in[a,b]$ (see [1]). Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, the assumptions follow from the observation that $\varepsilon=\varepsilon_{c}\circ(J\times\text{Id}_{[a,b]})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$. ∎
Proposition 3.18
Let $M$, $N$ and $L$ smooth manifolds modeled on a sequentially complete locally convex space such that $L$ admits a local addition and $p\in[1,\infty]$. If $f:L\times M\to N$ is a $C^{2}$-map and $\gamma\in AC_{L^{p}}([a,b],L)$ is fixed, then the map
$F:[a,b]\times M\to N,\quad(t,p)\mapsto f(\gamma(t),p)$
is continuous.
Proof.
It follows from the fact that $F=f\circ(\gamma\times\text{id}_{M})$. ∎
Proposition 3.19
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$. For each $q\in N$ the function $\zeta_{q}:[a,b]\to N$, $t\mapsto q$ is absolutely continuous and the map
$\zeta:N\to AC_{L^{p}}([a,b],N),\quad q\mapsto\zeta_{q}$
is smooth and a topological embedding.
Proof.
Consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 3.1. Let $(U,\varphi)$ be a chart around $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
If $x\in\varphi\left(U\cap\zeta^{-1}(\mathcal{U}_{\zeta_{q}})\right)$, then for each $t\in[a,b]$ we have
$\Psi_{\zeta_{q}}^{-1}\circ\zeta\circ\varphi^{-1}(x)(t)$ $=\theta_{N}^{-1}\left(\zeta_{q}(t),\zeta_{\varphi^{-1}(x)}(t)\right)$
$=\theta_{N}^{-1}\left(q,\varphi^{-1}(x)\right)$
$=\theta_{N}^{-1}\circ(q,\varphi^{-1}\circ\tilde{\zeta}_{x}(t))$
where $\tilde{\zeta}_{x}:[a,b]\to E$, $t\mapsto x$ is the constant function. Since the map
$\tilde{\zeta}:E\to AC_{L^{p}}([a,b],E),\quad x\mapsto\tilde{\zeta}_{x}$
is a continuous linear, setting the smooth map
$h:\varphi(U)\to TN,\quad z\to\theta_{N}^{-1}\circ(q,\varphi^{-1}(z))$
we have
$\Psi_{\zeta_{p}}^{-1}\circ\zeta\circ\varphi^{-1}=AC_{L^{p}}([a,b],h)\circ\tilde{\zeta}|_{\phi(U)}.$
Hence $\zeta$ is smooth. Moreover, if $t\in[a,b]$, then $\varepsilon_{t}\circ\zeta=\text{id}_{N}:N\to N$. ∎
Following more general cases of manifolds of mappings, such as the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition (see e.g. [3, 24]), we well study the tangent bundle of the manifold of absolute continuous functions.
Remark 3.20
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and let $TAC_{L^{p}}([a,b],N)$ be the tangent bundle of $AC_{L^{p}}([a,b],N)$. Since the point evaluation map $\varepsilon_{t}:AC_{L^{p}}([a,b],N)\to N$ is smooth for each $t\in[a,b]$, we have
$T\varepsilon_{t}:TAC_{L^{p}}([a,b],N)\to TN.$
For each $v\in TAC_{L^{p}}([a,b],N)$ we define the function
$\Theta_{N}(v):[a,b]\to TN,\quad\Theta_{N}(v)(t)=T\varepsilon_{t}(v).$
Proposition 3.21
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then $\Theta_{N}(v)\in{\Gamma_{AC}(\eta)}$ for each $v\in T_{\eta}AC_{L^{p}}([a,b],N)$ and the map
$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}(v)$
is an isomorphism of topological vector spaces.
Proof.
Let $\Sigma:\Omega\to N$ be a normalized local addition of $N$. Since ${\Gamma_{AC}(\eta)}$ is a vector space, we identify its tangent bundle with ${\Gamma_{AC}(\eta)}\times{\Gamma_{AC}(\eta)}$. Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
$T\Psi_{\eta}:T\mathcal{V}_{\eta}\simeq\mathcal{V}_{\eta}\times{\Gamma_{AC}(\eta)}\to TAC_{L^{p}}([a,b],N)$
is a diffeomorphism onto its image. Moreover,
$T\Psi_{\eta}:\{0\}\times{\Gamma_{AC}(\eta)}\to T_{\eta}AC_{L^{p}}([a,b],N)$
is an isomorphism of topological vector spaces. We will show that
$\Theta_{\eta}\circ T\Psi_{\eta}(0,\sigma)=\sigma$
for each $\sigma\in{\Gamma_{AC}(\eta)}$. Which is equivalent to show that
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)=\sigma(t)\quad\text{ for all }t\in[a,b].$
Working with the geometric point of view of tangent vectors, we see that $(0,\sigma)$ is equivalent to the curve $[s\mapsto s\sigma]$. Hence, for each $t\in[a,b]$ we have
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}\circ T\Psi_{\eta}([s\mapsto s\sigma])$
$=T\varepsilon_{t}([s\mapsto\Psi_{\eta}(s\sigma)])$
$=T\varepsilon_{t}([s\mapsto\Sigma(s\sigma)])$
$=[s\mapsto\Sigma|_{T_{\eta(t)}N}(s\sigma(t))]$
$=T_{0}\Sigma|_{T_{\eta(t)}N}([s\mapsto s\sigma(t)]).$
Since $\Sigma$ is normalized we have $T_{0}\Sigma|_{T_{\eta(t)}N}=\text{id}_{T_{\eta(t)}N}$ and
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)=\sigma(t).$
In consequence, for each $\sigma\in{\Gamma_{AC}(\eta)}$, there exists a $v\in T_{\eta}AC_{L^{p}}([a,b],N)$ with $v=T\Psi_{\eta}(0,\sigma)$ such that
$\Theta_{\eta}(v)=\sigma.$
Moreover, the function
$\Theta_{\eta}(v):[a,b]\to TN,\quad t\mapsto\Theta_{N}(v)(t)=\sigma(t)\in T_{\eta(t)}N$
is in $AC_{L^{p}}([a,b],TN)$ and verifies $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
Remark 3.22
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition and $p\in[1,\infty]$. Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, by Proposition 3.8, the map
$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is smooth. Moreover, for $\eta\in AC_{L^{p}}([a,b],N)$, we have
$AC_{L^{p}}([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
Then $AC_{L^{p}}([a,b],\pi_{TN})$ defines a vector bundle on $AC_{L^{p}}([a,b],N)$.
By Proposition 3.8, we can apply [3, Theorem A.12] for our particular case:
Proposition 3.23
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\pi_{TN}:TN\to N$ its tangent bundle. Then the map
$\Theta_{N}:TAC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],TN),\quad v\mapsto\Theta_{N}(v)$
is an isomorphism of vector bundles.
Proposition 3.24
Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a smooth map, then the tangent map of
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is given by
$TAC_{L^{p}}([a,b],f)=\Theta_{N}^{-1}\circ AC_{L^{p}}([a,b],Tf)\circ\Theta_{M}.$
Proof.
By Proposition 3.8, we know that $AC_{L^{p}}([a,b],f)$ is smooth, thus $TAC_{L^{p}}([a,b],f)$ makes sense. Let us consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
$T\Psi_{\eta}:\{0\}\times{\Gamma_{AC}(\eta)}\to T_{\eta}AC_{L^{p}}([a,b],M).$
For $t\in[a,b]$ we denote the point evaluation in $M$ and $N$ as $\varepsilon_{t}^{M}$ and $\varepsilon_{t}^{N}$ respectively, then for each $\sigma\in{\Gamma_{AC}(\eta)}$ we have
$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
$=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)([s\mapsto\Sigma_{M}(s\sigma)])$
$=T\varepsilon_{t}^{N}([s\mapsto f\circ\Sigma_{M}(s\sigma)])$
$=[s\mapsto\varepsilon_{t}\left(f\circ\Sigma_{M}(s\sigma)\right)]$
$=[s\mapsto f\circ\Sigma_{M}(s\sigma(t))]$
$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}M}([s\sigma(t)])$
$=Tf([s\mapsto s\sigma(t)])$
$=Tf(\sigma(t))$
$=AC_{L^{p}}([a,b],Tf)\circ T\varepsilon_{t}^{M}\circ T\Psi_{\eta}(0,\sigma).$
Hence
$\Theta_{N}\circ TAC_{L^{p}}([a,b],f)=AC_{L^{p}}([a,b],Tf)\circ\Theta_{M}.$
Example 3.25
Let $p\in[1,\infty]$. If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group (see [14, 33]). We will give an alternative proof of this.
Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the left action
$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v:=TL_{g}(v).$
If $\varphi:U\subseteq G\to V\subseteq T_{e}G$ is a chart in $e$ such that $\varphi(e)=0$, then the set
$\Omega_{\varphi}:=\bigcup\limits_{g\in G}g.V\subseteq TG$
and the map
$\Sigma_{\varphi}:\Omega_{\varphi}\to G,\quad v\mapsto\pi_{TG}(v)\left(\varphi^{-1}(\pi_{TG}(v)^{-1}.v)\right)$
defines a local addition for $G$ (see e.g. [23]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with the local adition $(\Omega_{\varphi},\Sigma_{\varphi})$.
Let $\mu_{G}:G\times G\to G$ and $\lambda_{G}:G\to G$ be the multiplication map and inversion maps on $G$ respectively, we define the multiplication map $\mu_{AC}$ and the inversion map $\lambda_{AC}$ on $AC_{L^{p}}([a,b],G)$ as
$\mu_{AC}:=AC_{L^{p}}([a,b],\mu_{G}):AC_{L^{p}}([a,b],G)\times AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G)$
and
$\lambda_{AC}:=AC_{L^{p}}([a,b],\lambda_{G}):AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G)$
that by Lemma 2.17 and Proposition 3.8 are smooth.
We observe that for the neutral element $\zeta_{e}:[a,b]\to G$, $t\mapsto e$ of $AC_{L^{p}}([a,b],G)$ we have
$\Gamma_{AC}(\zeta_{e})=AC_{L^{p}}([a,b],T_{e}G).$
If $\Psi_{\zeta_{e}}^{-1}:\mathcal{U}_{\zeta_{e}}\to\mathcal{V}_{\zeta_{e}}$ is a chart in ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$, then we have $\mathcal{U}_{\zeta_{e}}=AC([a,b],U)$ and ${\mathcal{V}_{\zeta_{e}}=AC([a,b],V)}$. Moreover, we see that
$\Psi_{\zeta_{e}}\circ AC_{L^{p}}([a,b],\varphi)(\eta)$ $=\Sigma_{\varphi}\circ(\varphi\circ\eta)$
$=\pi_{TG}(\varphi\circ\eta)\left(\varphi^{-1}(\pi_{TG}(\varphi\circ\eta)^{-1}.\varphi\circ\eta)\right)$
$=e\varphi^{-1}(e.\varphi\circ\eta)$
$=\eta.$
This enables us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ we have a chart given by
$AC_{L^{p}}([a,b],\varphi):AC_{L^{p}}([a,b],U)\to AC([a,b],V),\quad\eta\mapsto\varphi\circ\eta.$
4 Semiregularity of right half-Lie groups.
Definition 4.1
A group $G$, endowed with a smooth manifold structure modeled on a locally convex space, is called a right half-Lie group if it is a topological group and if for all $g\in G$, the right translations $\rho_{g}:G\to G,x\mapsto xg$ are smooth.
Remark 4.2
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space. We define the right action
$TG\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v)$ (4.1)
and consider its restriction
$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$ (4.2)
Unlike on Lie groups, on half-Lie groups the latter action may not be smooth.
The following proposition is direct application of Proposition 3.8.
Proposition 4.3
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],G)$ and $g\in G$ we define the function
$\eta.g(t):=\eta(t)g,\quad\text{for all }t\in[a,b].$
Then $\eta.g\in AC_{L^{p}}([a,b],G)$ for each $g\in G$ and the map
$AC_{L^{p}}([a,b],\rho_{g}):AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G),\quad\eta\mapsto\eta.g$
is smooth.
Remark 4.4
The smoothness of the map
$\mathcal{R}:AC_{L^{p}}([a,b],G)\times G\to AC_{L^{p}}([a,b],G),\quad(\eta,g)\mapsto\eta.g$
would imply the smoothness of the multiplication map on $G$. In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth (see [37]), the multiplication map on $G$ would be smooth as it coincides with the composition
$\varepsilon_{a}\circ\mathcal{R}\circ(\zeta,\text{id}_{G}):G\times G\to G,\quad(h,g)\mapsto hg.$
Definition 4.5
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],G)$. Let $\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolute continuity for $\eta$. For each $i\in\{1,...,n\}$ we denote
$\eta_{i}:=\varphi_{i}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
Then $\eta_{i}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{i}=[\gamma_{i}]$ with $\gamma_{i}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$. We define $\gamma:[a,b]\to TG$ via
$\gamma(t):=T(\varphi_{i})^{-1}(\eta_{i}(t),\gamma_{i}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }i\in\{1,...,n\},$
and $\gamma(b)=T(\varphi_{n})^{-1}(\eta_{n}(b),\gamma_{n}(b))$. Then $\gamma$ is measurable (see [14]) and we define
$\dot{\eta}:=[\gamma].$
Definition 4.6
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is $L^{p}$-semiregular if for each $[\gamma]\in L^{p}([0,1],T_{e}G)$, there exists an $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
${\dot{y}}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$ (4.3)
$y(0)$ $=e$ (4.4)
such that the differential equation satisfies local uniqueness of Carathéodory solutions in the sense of [18]. In this case, we define the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$ (4.5)
Additionally, if $G$ admits a local addition, we say that $G$ is $L^{p}$-regular if $G$ is $L^{p}$-semiregular and if the evolution map is smooth. The definition for the case of $L_{rc}^{\infty}$-semiregularity and $L_{rc}^{\infty}$-regularity is analogous.
Definition 4.7
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-semiregular if there is exists an open 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that for each $[\gamma]\in B$ there exists a $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
$\dot{y}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$
$y(0)$ $=e$
and the latter differential equation satisfies local uniqueness of solutions.
For our purpose, we will use the subdivision property [33, Lemma 2.17].
Lemma 4.8
Let $E$ be a locally convex space, $p\in[1,\infty]$ and $[\gamma]\in L^{p}([0,1],E)$. For each $n\in\mathbb{N}$ and $k\in\{0,1,...,n-1\}$ we define
$\gamma_{n,k}:[0,1]\to E,\quad\gamma_{n,k}(t):=\frac{1}{n}\gamma\left(\frac{k+t}{n}\right).$
Then $[\gamma_{n,k}]\in L^{p}([0,1],E)$. Moreover, for each $[\gamma]\in L^{p}([0,1],E)$ and continuous seminorm $q$ on $L^{p}([0,1],E)$, we have
$\sup_{k\in\{0,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
Lemma 4.9
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. Then $G$ is locally $L^{p}$-semiregular if and only if $G$ is $L^{p}$-semiregular.
Proof.
Let $G$ be locally $L^{p}$-semiregular. Then there exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta)\in AC_{L^{p}}([a,b],G)$ exists for each $[\beta]\in B$. Without loss of generality we assume that
$B=\{[\beta]\in L^{p}([0,1],T_{e}G):\lVert[\beta]\lVert_{L^{p},q}<1\}$
for some continuous seminorm $q$ on $L^{p}([0,1],T_{e}G)$.
Let $[\gamma]\in L^{p}([0,1],T_{e}G)$. By Lemma 4.8, we find $n\in\mathbb{N}$ such that
$[\gamma_{n,k}]\in B,\quad\text{ for all }k\in\{0,1,...,n-1\}.$
Since each map
$\alpha_{k}:L^{p}([0,1],T_{e}G)\to L^{p}([0,1],T_{e}G),\quad[\tau]\mapsto[\tau_{n,k}]$
is continuous linear, there is exists an open $\gamma$-neighborhood $W_{\gamma}$ such that
$\alpha_{k}(W_{\gamma})\subseteq B,\quad\text{for all }k\in\{0,1,...,n-1\}.$
For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we write $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
$\eta_{\beta}(t):=\eta_{0}(nt),\quad\text{ if }t\in[0,1/n]$
and
$\eta_{\beta}(t):=\eta_{k}(nt-k).\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)},\quad\text{ if }t\in[k/n,(k+1)/n]\text{ with }k\in\{1,..,n-1\}.$
Then the function $\eta_{\beta}$ is continuous and by Proposition 4.3 we have
$\eta_{\beta}|_{[k/n,(k+1)/n]}\in AC_{L^{p}}\left([k/n,(k+1)/n],G\right).$
Thus $\eta_{\beta}\in AC_{L^{p}}([a,b],G)$. If $t\in[0,\frac{1}{n}]$ we have $\eta_{\beta}(0)=e$ and
$\dot{\eta_{\beta}}(t)$ $=\dot{\eta_{0}}(nt)$
$=n\beta_{n,0}(nt).\eta_{0}(nt)$
$=n\frac{1}{n}\beta\left(\frac{0+nt}{n}\right).\eta_{0}(nt)$
$=\beta(t).\eta_{\beta}(t).$
For $k\in\{1,..,n-1\}$ we have
$\eta_{\beta}\left(\frac{k}{n}\right)=\eta_{k}\left(n\frac{k}{n}-k\right).\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}=\eta_{k-1}(1)....\eta_{0}(1)$
and for $t\in[\frac{k}{n},\frac{k+1}{n}]$
$\dot{\eta_{\beta}}(t)$ $=n\beta_{n,k}(nt-k)\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}.\eta_{k}(nt-k)$
$=\beta(t).\eta_{k}(nt-k)\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}$
$=\beta(t).\eta_{\beta}(t).$
Thus $\eta_{\beta}=\operatorname{Evol}(\beta)$ and in particular, $\eta_{\gamma}=\operatorname{Evol}(\gamma)$. Hence $G$ is $L^{p}$-semiregular.
The reciprocal is trivial. ∎
Definition 4.10
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
The following lemma is just an application of [14, Proposition 5.25] or [32, Proposition 4.11] to our case.
Lemma 4.11
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admit a local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G)\to C([0,1],G),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma):=\eta_{\gamma}.$
Then, $\operatorname{Evol}_{C}$ is continuous if and only if there exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that the restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
Proof.
Since $G$ is a topological group, the map
$C([0,1],G)\times C([0,1],G)\to C([0,1],G),\quad(\eta,\xi)\mapsto\eta.(\xi(1))$
is continuous. Following Lemma 4.9, if $[\gamma]\in L^{p}([0,1],T_{e}G)$, for each $[\beta]\in W_{\gamma}$, we have that the construction $\operatorname{Evol}_{C}(\beta)$ implies that $\left(\text{Evol}_{C}\right)|_{W_{\gamma}}$ is just the product of composition of continuous maps. ∎
Lemma 4.12
Let $1\leq p<\infty$, $E_{1}$ be a Fréchet space, $E_{2}$ and $F$ be locally convex spaces and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
$\tilde{f}:L^{p}([0,1],E_{1})\times C([0,1],U)\to L^{p}([0,1],F),\quad([\gamma],\eta)\mapsto f\circ([\gamma],\eta)$
is $C^{r}$.
Proof.
Let $(\gamma_{0},\eta_{0})\in L^{p}([0,1],E_{1})\times C([0,1],U)$ and $\varepsilon>0$. Let $\beta:F\to[0,\infty)$ be a continuous seminorm and $K:=\eta_{0}([0,1])\subseteq U$. Since $f$ is continuous and linear in the first argument, for $y\in K$ there exists a $y$-neighborhood $V_{y}\subseteq U$ and a continuous seminorm $\kappa_{y}:E_{1}\to[0,\infty)$ such that
$f\left(B_{1}^{\kappa_{y}}(0)\times V_{y}\right)\subseteq B_{1}^{\beta}(0).$
By compactness of $K$, there exists a finite numbers of $y_{1},...,y_{n}\in K$ such that
$K\subseteq V:=\cup_{i=1}^{n}V_{y_{i}}.$
We define the continuous seminorm $\kappa:E_{1}\to[0,\infty)$ by $\kappa:=\kappa_{y_{1}}+...+\kappa_{y_{n}}$ and we have
$\beta\big{(}f(x,b)\big{)}\leq\kappa(x),\quad\text{for all }x\in E_{1},b\in V$
and $C([0,1],V)$ is open in $C([0,1],U)$. We estimate
[…]
$\leq\lVert\gamma_{0}\lVert_{L^{p},\kappa}.$
Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$.
Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$ (see [14, Lemma A.4]) there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5.$
Additionally, since the maps
$C([0,1],U)\to C([0,1],F),\quad\eta\mapsto f\circ(\gamma_{c},\eta)$
and the inclusion map $C([0,1],F)\to L^{p}([0,1],F)$ are continuous, we have that
$C([0,1],V)\to L^{p}([0,1],F),\quad\eta\mapsto f\circ(\gamma_{c},\eta)$
is continuous. Hence there exists an open neighborhood $W$ of $\eta_{0}\in C([0,1],V)$ such that
$\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5,\quad\text{for all }\eta\in W.$
With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we estimate
$\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$ $\leq\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$=\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta_{0})-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$\leq\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$\leq\Big{\lVert}\gamma-\gamma_{c}\Big{\lVert}_{L^{p},\kappa}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\Big{\lVert}\gamma_{c}-\gamma_{0}\Big{\lVert}_{L^{p},\kappa}.$
Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we have
$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5.$
Thus
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+\varepsilon/5=\varepsilon.$
Thus $\tilde{f}$ is continuous. By linearity in the first variable, the map $\overline{f}$ has a continuous differential in the first variable
$d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F).$
Let consider the map $f$ as a $C^{1}$-map. If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
[…]
whenever $y+[0,1]ty_{1}\subseteq U$. Given that the map $d_{2}f:E_{1}\times U\times E_{2}\to F$ is continuous, identifying $C([0,1],E_{1}\times E_{2})$ with $C([0,1],E_{1})\times C([0,1],E_{2})$, we have that the map
$\widetilde{d_{2}f}:L^{p}([0,1],E_{1})\times C([0,1],U)\times C([0,1],E_{2})\to L^{p}([0,1],F),\quad(\gamma,\eta,\eta_{1})\mapsto d_{2}f(\gamma,\eta,\eta_{1})$
is continuous. For $t\in[0,1]$ fixed, we denote
$g_{t}:C([0,1],F)\to F,\quad\gamma\mapsto\gamma(t).$
Since the family of maps $\{g_{t}:t\in[0,1]\}$ separate points on $C([0,1],F)$, we have that the equality
[…]
is valid for each $\gamma_{c}\in C([0,1],E_{1})$, $\eta\in C([0,1],U)$, $\eta_{1}\in C([0,1],E_{2})$ and $t\in\mathbb{R}^{\times}$ such that
$\eta+[0,1]t\eta_{1}\in C([0,1],U).$
By density of $C([0,1],E_{1})$ on $L^{p}([0,1],E_{1})$ and continuity of $\tilde{f}$, for $\gamma\in L^{p}([0,1],E)$ the equation verifies
[…]
Let $\gamma$, $\eta$ and $\eta_{0}$ be fixed, then the map $(t,s)\mapsto\widetilde{d_{2}f}(\gamma,\eta+ts\eta_{1},\eta_{1})$ is continuous, including in $t=0$. Then, taken the limit $t\to 0$ in the equality we obtain
$d_{2}\tilde{f}(\gamma,\eta,\eta_{1})=\widetilde{d_{2}f}(\gamma,\eta,\eta_{1}).$
Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\widetilde{f}$ is $C^{1}$. Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
$d\tilde{f}(\gamma,\eta,\gamma_{1},\eta_{1})$ $=d_{1}\tilde{f}(\gamma,\gamma_{1},\eta)+d_{2}\tilde{f}(\gamma,\eta,\eta_{1})$
$=\tilde{f}(\gamma_{1},\eta)+\widetilde{d_{2}f}(\gamma,\eta,\eta_{1}).$
By the induction hypothesis $\tilde{f}$ and $d_{2}f$ are $C^{r-1}$, hence $\widetilde{d_{2}f}$ is $C^{r-1}$ with $\widetilde{d_{2}f}=d_{2}\tilde{f}$, thus $\tilde{f}$ is $C^{r}$. ∎
For the case $L_{rc}^{\infty}$ we recall [14, Proposition 2.3].
Lemma 4.13
Let $E_{1},E_{2}$ and $F$ be integral complete locally convex spaces and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
$\tilde{f}:L_{rc}^{\infty}([0,1],E_{1})\times C([0,1],U)\to L_{rc}^{\infty}([0,1],F),\quad([\gamma],\eta)\mapsto f\circ([\gamma],\eta)$
is $C^{r}$.
Theorem 4.14
Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits a local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G)\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
If the restriction of the right action
$\tau:T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g$
is continuous, then the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
Proof.
Let $[\gamma]\in L^{p}([0,1],T_{e}G)$. Let $\mathcal{P}_{[\gamma]}$ be a open neighborhood of $\operatorname{Evol}_{C}(\gamma)$ in $C([0,1],G)$, then there exists a partition $P=\{t_{0},...,t_{n}\}$ of $[0,1]$ and a family of charts $\varphi_{i}:U_{i}\to V_{i}$ such that
$\mathcal{P}_{[\gamma]}=\bigcap_{i=1}^{n}\{\eta\in C([t_{i-1},t_{i}],U_{i}):\eta([t_{i-1},t_{i}])\subseteq U_{i}\}.$
Let $\mathcal{Q}_{[\gamma]}=\operatorname{Evol}_{C}^{-1}(\mathcal{P}_{[\gamma]})$, then $\mathcal{Q}_{[\gamma]}$ is an open neighborhood of ${[\gamma]}$ in $L^{p}([0,1],T_{e}G)$. We will show that the map
$\operatorname{Evol}|_{\mathcal{Q}_{[\gamma]}}:\mathcal{Q}_{[\gamma]}\to\mathcal{P}_{[\gamma]}\cap AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. The map
$\Phi_{1}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G),\quad[\beta]\mapsto\left([\beta|_{[t_{i-1},t_{i}]}]\right)_{i=1}^{n}$
is a topological embedding. For each $i\in\{1,...,n\}$ we have
$\operatorname{Evol}_{C}(\beta|_{[t_{i-1},t_{i}]})=\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]},\quad\text{ for all }[\beta]\in\mathcal{Q}_{[\gamma]}$
and each map $C([t_{i-1},t_{i}],U_{i})\to C([t_{i-1},t_{i}],V_{i})$, $\eta\mapsto\varphi_{i}\circ\eta$ is a homeomorphism. This allow us to define the continuous map
$\Phi_{2}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}C([t_{i-1},t_{i}],V_{i}),\quad[\beta]\mapsto\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
We consider the map
$f_{i}:T_{e}G\times V_{i}\to E,\quad(v,x)\mapsto d\varphi_{i}\circ\tau(v,\varphi_{i}^{-1}(x)).$
Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function
$f_{i}(\cdot,x):T_{e}G\to E,\quad v\mapsto f_{i}(v,x)$
is linear. Hence by Lemma 4.12 the map
$\tilde{f}_{i}:L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to L^{p}([t_{i-1},t_{i}],E),\quad([\beta],\eta)\mapsto f_{i}\circ([\beta],\eta)$
is continuous. We denote
$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i})\right)_{i=1}^{n}.$
Then $F$ is continuous, where for each $i\in\{1,...,n\}$ we have
$f_{i}\circ([\beta_{i}],\eta_{i})=d\varphi_{i}\circ\tau([\beta_{i}],\varphi_{i}^{-1}\circ\eta_{i}).$
This allow us to define the continuous map
$\Phi_{3}:\mathcal{Q}_{[\gamma]}$ $\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E)\times C([t_{i-1},t_{i}],V_{i}),$
$\beta$ $\mapsto\Big{(}f_{i}\circ\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right),\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}_{i=1}^{n}$
where for each $i\in\{1,...,n\}$ we have
$f_{i}\circ\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}\circ\operatorname{Evol}_{C}|_{[t_{i-1},t_{i}]}\right)$ $=d\varphi_{i}\circ\tau\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}^{-1}\circ\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)$
$=d\varphi_{i}\circ\tau\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)$
$=d\varphi_{i}\circ\left(\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime}$
$=\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime}.$
Hence
$\Phi_{3}([\beta])=\Big{(}\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime},\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}_{i=1}^{n},\quad\text{ for all }[\beta]\in\mathcal{Q}_{[\gamma]}.$
We set the topological embedding (see Lemma 2.13)
$\Psi_{i}:AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to L^{p}([t_{i-1},t_{i}],V_{i})\times C([t_{i-1},t_{i}],V_{i}),\quad\alpha\mapsto(\alpha^{\prime},\alpha)$
and
$\Psi:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i}),\quad(\alpha_{i})_{i=1}^{n}\mapsto(\alpha_{i}^{\prime},\alpha_{i})_{i=1}^{n}.$
Thus $\text{Im}(\Phi_{3})\subseteq\text{Im}(\Psi)$. For each $i\in\{1,...,n\}$ we note that
$\left(\Psi_{i}|^{\text{Im}(\Psi_{i})}\right)^{-1}\Big{(}(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]})^{\prime},\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}=\varphi_{i}\circ\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}$
for all $[\beta]\in\mathcal{Q}_{[\gamma]}$. We set the continuous map
$\Phi_{4}:=\left(\Psi|^{\text{Im}(\Psi)}\right)^{-1}\circ\Phi_{3}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i}),\quad[\beta]\mapsto\left(\varphi_{i}\circ\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
Let denote map
$\Phi_{5}:AC_{L^{p}}([0,1],G)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],G),\quad\eta\mapsto(\eta|_{[t_{i-1},t_{i}]})_{i=1}^{n}.$
which is a homeomorphism onto its image the homeomorphism onto its image (see [37, Proposition 4.13]) and the homeomorphism
$\Phi_{6}:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],U_{i}),\quad\tau_{i}\mapsto(\varphi_{i}^{-1}\circ\tau_{i})_{i=1}^{n}.$
We see that
$\Phi_{6}\circ\Phi_{4}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],U_{i}),\quad[\beta]\mapsto\left(\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
Since each function $\operatorname{Evol}(\beta)$ is continuous, we have that $(\Phi_{6}\circ\Phi_{4})\left(\mathcal{Q}_{[\gamma]}\right)\subseteq\text{Im}(\Phi_{5})$, hence
$\operatorname{Evol}|_{\mathcal{Q}_{[\gamma]}}=\left(\Phi_{5}|^{\text{Im}(\Phi_{5})}\right)^{-1}\circ\Phi_{6}\circ\Phi_{4}.$
Thus the evolution map $\operatorname{Evol}$ is continuous.
For the case $L_{rc}^{\infty}$, by Lemma 4.13, the proof is analogous since the map
$\tilde{f}_{i}:L_{rc}^{\infty}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to L_{rc}^{\infty}([t_{i-1},t_{i}],E),\quad([\beta],\eta)\mapsto f_{i}\circ([\beta],\eta)$
is continous for each $i\in\{1,...,n\}$. ∎
5 Semiregularity of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$
Let $n,m,r\in\mathbb{N}$. We consider the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
$\lVert\cdot\lVert_{L}:C(\mathbb{R}^{n},\mathbb{R}^{m})\to[0,\infty),\quad\phi\mapsto\sup_{x\in L}\lvert\phi(x)\lvert$
for each non empty compact subset $L\subset\mathbb{R}^{n}$. Let $C^{r}(\mathbb{R}^{n},\mathbb{R}^{m})$ be the Fréchet space of all $C^{r}$-maps $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$, whose topology is the compact-open $C^{r}$-topology, i.e., the initial topology with respect to the maps
$C^{r}(\mathbb{R}^{n},\mathbb{R}^{m})\to C(\mathbb{R}^{n},\mathbb{R}^{m}),\quad\phi\mapsto\frac{\partial^{\alpha}\phi}{\partial x^{\alpha}}$
for $\alpha\in\mathbb{N}_{0}^{n}$ with $\lvert\alpha\lvert\leq r$. Let $K\subseteq\mathbb{R}^{n}$ be a non empty compact subset, we define the Banach space of $C^{r}$-maps supported in $K$ as
$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$ (5.1)
endowed with the induced topology.
We denote derivative map
$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$ (5.2)
where $f^{\prime}(x)$ is the Jacobian matrix of $f$.
Definition 5.1
Let $\text{Diff}^{r}(\mathbb{R}^{n})$ be the set of $C^{r}$-diffeomorphisms $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The set $\text{Diff}^{r}(\mathbb{R}^{n})$ is a group under the composition and we define the subgroup $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ of $C^{r}$-diffeomorphisms which are supported in a compact set $K\subset\mathbb{R}^{n}$ as
$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$ (5.3)
Let $\mathcal{V}_{K}:=\{\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}):\phi\in\text{Diff}_{K}^{r}(\mathbb{R}^{n})\}$, then $\mathcal{V}_{K}$ is open in $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ and the map
$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$ (5.4)
is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [16, Proposition 14.6]). On the set $\mathcal{V}_{K}$ we define a group multiplication by
$\phi\ast\psi:=\Phi\left(\Phi^{-1}(\phi)\circ\Phi^{-1}(\psi)\right)=\psi+\phi\circ(\text{id}_{\mathbb{R}^{n}}+\psi),\quad\text{ for each }\phi,\psi\in\mathcal{V}_{K}$
with the constant function 00 as neutral element.
Remark 5.2
We will study the $L^{1}$-semiregularity of $\mathcal{V}_{K}$ instead of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$. Since $\mathcal{V}_{K}$ is an open set of a locally convex space, we have
$T\mathcal{V}_{K}=\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}).$
For $\psi\in\mathcal{V}_{K}$ fixed, we have the right translation
$\rho_{\psi}:\mathcal{V}_{K}\to\mathcal{V}_{K},\quad\phi\mapsto\psi+\phi\circ(\text{id}_{\mathbb{R}^{n}}+\psi)$
and its derivative
$d\rho_{\psi}:\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}),\quad(\phi,\varphi)\mapsto\varphi\circ(\text{id}_{\mathbb{R}^{n}}+\psi).$
Identifying $T_{0}\mathcal{V}_{K}$ with $\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, the restricted right action is given by
($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$ (5.5)
Hence the right action is continuous (see e.g. [15]).
If $(\phi,\varphi)\in\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, we have
$(\phi,\varphi).\psi:=T\rho_{\psi}(\phi,\varphi)=(\phi\ast\psi,\varphi\circ(\text{id}_{\mathbb{R}^{n}}+\psi)).$
If $(0,\gamma)\in\mathcal{L}^{1}([0,1],T_{0}\mathcal{V}_{K})$, we want to find a function $\eta\in AC_{L^{p}}([0,1],\mathcal{V}_{K})$ such that $\eta^{\prime}=\gamma.\eta$, i.e.
$(\eta(t),\eta^{\prime}(t))=\Big{(}\eta(t),\gamma(t)\circ(\text{id}_{\mathbb{R}^{n}}+\eta(t))\Big{)},\text{ for almost all }t\in[0,1].$
In other words, for $\gamma\in\mathcal{L}^{1}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$, we need a function $\eta\in AC_{L^{p}}([0,1],\mathcal{V}_{K})$ such that $\eta(0)=0$ and
[…]
Setting $\zeta=\text{id}_{\mathbb{R}^{n}}+\eta(s)$, this is equivalent to
[…]
Let $\varepsilon_{x}:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to\mathbb{R}^{n}$ be the point evaluation map for $x\in\mathbb{R}^{n}$, then each $\varepsilon_{x}$ is continuous linear and the family of maps $(\varepsilon_{x})_{x\in\mathbb{R}^{n}}$ separate points on $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, hence the equation holds if and only if the functions $\zeta_{x}:=\varepsilon_{x}\circ\zeta\in AC_{L^{1}}([0,1],\mathbb{R}^{n})$ satisfy
[…]
Theorem 5.3
Let $1\leq p<\infty$. If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L^{p}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad\gamma\mapsto\eta_{\gamma}.$
is continuous.
Proof.
Following the discussion in Remark 5.2, we will show that $\mathcal{V}_{K}$ is locally $L^{p}$-semiregular.
We define the continuous seminorm
$\alpha:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to[0,\infty),\quad\alpha(f):=\lVert f\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}=\sup_{x\in\mathbb{R}^{n}}\lVert Df(x)\lVert_{op}.$
For $10<L<$ 1, we denote the open ball centered in $0\in L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ by
$B_{L}:=\left\{[\gamma]\in L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})):\lVert\gamma\lVert_{L^{p},\alpha}:=\left(\int_{0}^{1}(\alpha\circ\gamma)^{p}(t)dt\right)^{1/p}<L\right\}.$
For $x\in\mathbb{R}^{n}$, we define the smooth map
$c:\mathbb{R}^{n}\to C([0,1],\mathbb{R}^{n}),\quad x\mapsto[t\mapsto x].$
Let $J:L^{1}([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ be the continuous linear operator
[…]
Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{\infty,r}$ in the sense of [1] and hence $C^{r}$. By Lemma 4.12 the map
$\Phi_{n}:L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\times C([0,1],\mathbb{R}^{n})\to L^{p}([0,1],\mathbb{R}^{n}),\quad\Phi_{n}([\gamma],\zeta):=\varepsilon\circ([\gamma],\zeta)$
is well defined and is $C^{r}$. We define the operator $T:B_{L}\times\mathbb{R}^{n}\times C([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ via
$T([\gamma],x,\zeta)(t)$ $:=c_{x}(t)+J(\Phi_{n}([\gamma],\zeta))(t)$
[…]
for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$ in the sense of [2].
Let $[\gamma]\in B_{L}$ and $x\in\mathbb{R}^{n}$ be fixed, for $\zeta_{1},\zeta_{2}\in C([0,1],\mathbb{R}^{n})$, by the Mean Value Theorem we have
$\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)=\int_{0}^{1}\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))ds.$
Thus
$\lVert T([\gamma],x,\zeta_{2})-T([\gamma],x,\zeta_{1})\lVert_{\infty}$ $=\sup_{t\in[0,1]}\left\lVert\int_{0}^{t}(\gamma(s)\circ\zeta_{2}(s)-\gamma(s)\circ\zeta_{1}(s))ds\right\lVert_{\infty}$
$\leq\sup_{t\in[0,1]}\int_{0}^{1}\left\lVert\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))\right\lVert_{\infty}ds$
$\leq\sup_{t\in[0,1]}\int_{0}^{1}\lVert D\gamma(t)\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}dt$
$=\sup_{t\in[0,1]}\int_{0}^{1}\alpha(\gamma(t))dt\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq\lVert\gamma\lVert_{L^{1},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq\lVert\gamma\lVert_{L^{p},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq L\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}.$
Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and an $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
$T([\gamma],x,\zeta_{[\gamma],x})=\zeta_{[\gamma],x}.$
By [14, Lemma 6.2] the map
$F_{1}:B_{L}\times\mathbb{R}^{n}\to C([0,1],\mathbb{R}^{n}),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
is $C^{r}$. By the exponential laws, we have:
  • a)
    $F_{2}:\left(B_{L}\times\mathbb{R}^{n}\right)\times[0,1]\to\mathbb{R}^{n},\quad([\gamma],x,t)\mapsto\zeta_{[\gamma],x}(t)$ is $C^{r,0}$.
  • b)
    $F_{3}:B_{L}\times[0,1]\to C^{r}(\mathbb{R}^{n},\mathbb{R}^{n}),\quad([\gamma],t)\mapsto\zeta_{[\gamma],\cdot}(t)$ is $C^{r,0}$.
  • c)
    $F_{4}:B_{L}\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto\zeta_{[\gamma]}$ is continuous.
Let $x\in\mathbb{R}^{n}\setminus K$ and $\zeta_{[\gamma,x}$ be the fixed point of
[…]
Since $\gamma(t)(x)=0$ for each $t\in[0,1]$, the constant map $c_{x}$ is also a fixed point and by uniqueness of solutions we have $\zeta_{[\gamma],x}=c_{x}$. Let $\text{I}:[0,1]\to C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, $t\mapsto\text{id}_{\mathbb{R}^{n}}$ the constant map mapping to the identity map. We define
$S:B_{L}\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
Then for each $t\in[0,1]$ we have $S([\gamma])(t)\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$. Moreover, since $S$ is continuous and $\mathcal{V}_{K}$ is open, there is exists an 0-neighborhood $B\subseteq B_{L}$ such that $S(B)\subseteq\mathcal{V}_{K}$. Hence, we can consider the map $S$ as the continuous map
$\tilde{S}:B\to C([0,1],\mathcal{V}_{K}),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I},$
where, by the discussion in Remark 5.2, if $[\gamma]\in B$ we have that $\eta:=\tilde{S}([\gamma])$ is solution of
[…]
Hence $\operatorname{Evol}(\gamma)=\eta$. In consequence, $\mathcal{V_{K}}$ is locally $L^{p}$-semiregular and thus, by Lemma 4.9, the half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L^{p}$-semiregular.
Moreover, the evolution map restricted to the 00-neighborhood $B$ is given by
$\operatorname{Evol}|_{B}:B\subseteq L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\tilde{S}(\gamma).$
Since $\tilde{S}$ is continuous, by Lemma 4.11 the evolution map with continuous values $\operatorname{Evol}_{C}$ is continuous. Moreover, since the restricted right action of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is continuous, by Theorem 4.14, the evolution map
$\operatorname{Evol}:L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
Proceeding exactly as the case $L^{p}$, for the case $L_{rc}^{\infty}$ we have the same result.
Proposition 5.4
If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L_{rc}^{\infty}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L_{rc}^{\infty}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))\to AC_{L_{rc}^{\infty}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad\gamma\mapsto\eta_{\gamma}$
is continuous.
Proof.
For $10<L<$ 1, we define the open ball centered in $0\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ via
$B_{L}:=\{[\gamma]\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})):\lVert\gamma\lVert_{L_{rc}^{\infty},\alpha}:=\operatorname*{ess\,sup}_{t\in[0,1]}\alpha(\gamma(t))<L\}$
Following the same notation as Proposition 1.7, since the evaluation map $\varepsilon$ is $C^{r,\infty}$, by Lemma 4.13 the map
$\Phi_{n}:L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\times C([0,1],\mathbb{R}^{n})\to L_{rc}^{\infty}([0,1],\mathbb{R}^{n}),\quad\Phi_{n}([\gamma],\zeta):=\varepsilon\circ([\gamma,\zeta])$
is $C^{r}$. Hence the operator $T:B_{L}\times\mathbb{R}^{n}\times C([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ given by
[…]
is a $C^{\infty,\infty,r}$-map. Let $[\gamma]\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ and $x\in\mathbb{R}^{n}$ be fixed, for $\zeta_{1},\zeta_{2}\in C([0,1],\mathbb{R}^{n})$, by the Mean Value Theorem we have
$\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)=\int_{0}^{1}\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))ds$
and
$\lVert\Phi_{n}([\gamma],\zeta_{2})(t)-\Phi_{n}([\gamma],\zeta_{1})(t)\lVert_{\infty}$ $=\lVert\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)\lVert_{\infty}$
$\leq\int_{0}^{1}\lVert\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))\lVert_{\infty}ds$
$\leq\lVert D\gamma(t)\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}$
$=\alpha(\gamma(t))\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}.$
Hence
$\lVert\Phi_{n}([\gamma],\zeta_{2})-\Phi_{n}([\gamma],\zeta_{1})\lVert_{L_{rc}^{\infty},\infty}\leq\lVert\gamma\lVert_{L_{rc}^{\infty},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}.$
Thus $T([\gamma],x,\cdot)$ is a $C^{r}$-map and a $L$-contraction. With this, following the same steps as Proposition 1.7 we can show that $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L_{rc}^{\infty}$-semiregular and there exists a 00-neighborhood $B$ in $L_{rc}^{\infty}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))$ such that the restricted map evolution map
$\operatorname{Evol}_{C}|_{B}:B\to AC_{L_{rc}^{\infty}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma)$
is continuous. By Lemma 4.11 the evolution map with continuous values $\operatorname{Evol}_{C}$ is continuous. Moreover, since the restricted right action of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is continuous, by Theorem 4.14, the evolution map
$\operatorname{Evol}:L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L_{rc}^{\infty}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
6 $L^{p}$-Semiregularity of $\text{Diff}^{\text{ }r}(M)$
Let $r\in\mathbb{N}$. Let $M$ be a compact smooth manifold without boundary, $g$ be a smooth Riemannian metric on $M$ and $\exp:\mathcal{D}\subseteq TM\to M$ be the Riemannian exponential map, with $\exp_{p}:=\exp|_{T_{p}M}$ for each $p\in M$. Let $\pi_{TM}:TM\to M$ be the bundle projection and $\mathcal{W}\subseteq\mathcal{D}$ be an open neighborhood of the zero-section such that the map
$(\pi_{TM},\exp):\mathcal{W}\subseteq TM\to M\times M,\quad v\mapsto(\pi_{TM}(v),\exp(v))$
is a diffeomorphism into its image. In particular, if $\mathcal{W}_{p}:=\mathcal{W}\cap T_{p}M$, the map
$\exp_{p}|_{\mathcal{W}_{p}}:\mathcal{W}_{p}\subseteq T_{p}M\to\exp_{p}\left(\mathcal{W}_{p}\right)\subseteq M$
is a $C^{\infty}$-diffeomorphism for each $p\in M$.
Let $\text{Diff}^{\text{ }r}(M)$ be the set of all $C^{r}$-diffeomorphism $\phi:M\to M$. For each $\phi\in\text{Diff}^{\text{ }r}(M)$ we define the Banach space of $C^{r}$-sections
$\Gamma_{C^{r}}(\phi):=\{X\in C^{r}(M,TM):\pi_{TM}\circ X=\phi\}$
and the open 00-neighborhood
$\mathcal{V}_{\phi}:=\{X\in\Gamma_{C^{r}}(\phi):X(M)\subseteq\mathcal{W}\}.$
If
$\mathcal{U}_{\phi}:=\{\exp\circ X\in\text{Diff}^{\text{ }r}(M):X\in\mathcal{V}_{\phi}\}$
then the map
$\Psi_{\phi}:\mathcal{U}_{\phi}\to\mathcal{V}_{\phi},\quad\Psi(\phi)(p):=\left(\exp_{p}|_{\mathcal{W}_{p}}\right)^{-1}(\phi(p))$
define a chart for $\text{Diff}^{\text{ }r}(M)$ on $\phi$, with inverse given by
$\Psi_{\phi}^{-1}:\mathcal{V}_{\phi}\to\mathcal{U}_{\phi},\quad X\mapsto\exp\circ X,$
turning $\text{Diff}^{\text{ }r}(M)$ into a Banach manifold. Moreover, under the composition the manifold $\text{Diff}^{\text{ }r}(M)$ becomes a right half-Lie group (see e.g. [16]). We see that for the neutral element $e=\text{id}_{M}$, the tangent space $T_{e}\text{Diff}^{\text{ }r}(M)$ coincide with space of $C^{r}$-vector fields $X:M\to TM$, denoted by $\mathcal{X}^{r}(M)$.
The restricted right action
$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$ (6.1)
is continuous (see e.g. [15]).
Let $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, we will study the equation
$\dot{\eta}(t)$ $=\gamma(t).\eta(t),\quad t\in[0,1]$ (6.2)
$\eta(0)$ $=\text{id}_{M}.$ (6.3)
Definition 6.1
Let $(E,\lVert\cdot\lVert_{E})$ be a normed space and $U\subseteq E$ be a subset. We say that a function $f:[a,b]\times U\to E$ satisfies an $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b],$
where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constants for $f(t,\cdot):U\to E$.
Definition 6.2
Let $M$ be a $C^{1}$-manifold modeled on a normed space $E$, $J\subseteq\mathbb{R}$ be a non-degenerate interval and $f:J\times M\to TM$ be a function with $f(t,p)\in T_{p}M$ for all $(t,p)\in J\times M$. We say that $f$ satisfies a local $L^{1}$-Lipschitz condition if for all $t_{0}\in J$ and $P\in M$, there exists a chart $\kappa:U_{\kappa}\to V_{\kappa}$ of $M$ on $p$ and a relatively open subinterval $[a,b]\subseteq J$ which is a neighborhood of $t_{0}$ in $J$ such that the map
$f_{\kappa}:[a,b]\times V_{\kappa}\to E,\quad(t,x)\mapsto d\kappa\left(f(t,\kappa^{-1}(y))\right)$
satisfies an $L^{1}$-Lipschitz condition.
Remark 6.3
Let $(E,\lVert\cdot\lVert)$ be a normed space and $f:[a,b]\times M\to TM$ be a map with $f(t,q)\in T_{q}M$ for all $(t,q)\in[a,b]\times M$, which satisfies a local $L^{1}$-condition. If $\tau,\eta:[a,b]\to M$ are two $AC_{L^{1}}$-Carathéodory solutions to
$y^{\prime}=f(t,y)$
satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [14, Proposition 10.5]. If the initial value problem
$y^{\prime}(t)$ $=f(t,y(t)),\quad t\in[a,b]$
$y(0)$ $=q$
has a (necessarily unique) solution $\gamma_{q}:[a,b]\to M$ for each $q\in M$, then we say that $f$ admits a global flow for initial time $t_{0}$ and write
$F_{t,a}^{f}(q):=\gamma_{q}(t),$
for all $t\in[a,b]$ and $q\in M$.
The following result can be found in the appendix of [14] for the context of $AC_{L^{1}}$-functions with values in the Lie group of $C^{\infty}$-diffeomorphism with compact support $\text{Diff}_{c}^{\infty}(M)$. However, this result is also valid in our context.
Theorem 6.4
Let $1\leq p<\infty$ and $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, then map
$\overline{\gamma}:[0,1]\times M\to TM,\quad\overline{\gamma}(t,q):=\gamma(t)(q)$
satisfies a local $L^{1}$-Lipschitz condition. Let $\eta\in AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M))$ with $\eta(0)=\text{id}_{M}$. Then $\eta=\operatorname{Evol}([\gamma])$ if and only if $f$ admits global flow $F$ for initial time $t_{0}=0$ and
$\eta(t)(p)=F_{t,0}^{\overline{\gamma}}(q),$
for all $t\in[0,1]$ and $q\in M$.
The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [14, Section 11]. We recommend to the reader to read the reference to see some steps in detail.
Theorem 6.5
Let $M$ be a compact smooth manifold and $1\leq p<\infty$. If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}\text{Diff}^{\text{ }r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\eta_{[\gamma]}$
is continuous.
Proof.
For $\phi\in\text{Diff}^{\text{ }r}(M)$, we can choose a finite locally finite cover $(U_{i})_{i=1}^{n}$ of $M$ by relatively compact, open subsets sets and charts $\kappa_{i}:U_{i}\to B_{5}(0)\subseteq\mathbb{R}^{n}$ such that the family of open sets $\kappa_{i}^{-1}(B_{1}(0))$ cover $M$ and $U_{i}\subseteq\phi^{-1}(U_{\psi_{i}})$ for some chart $\psi_{i}:U_{\psi_{i}}\to V_{\psi_{i}}$ of $M$ (see e.g. [25]). For $X\in\Gamma_{C^{r}}(\phi)$ we write
$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{m}.$
Then, for each $\ell\in[1,5]$ the map
$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{m}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
is a linear topological embedding with closed image (see [3]).
Let $\phi=\text{id}_{M}$, then $\Gamma_{C^{r}}(\text{id}_{M})=\mathcal{X}^{r}(M)$. Doing the corresponding identifications of product spaces, we have the linear topological embeddings with closed image
$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
and
$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
Before we study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the topological embedding $R_{1}$ allows us to study it first for
$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})).$
We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{m})\to[0,\infty)$ given by
$q(\phi):=\sup_{x\in\overline{B}_{4}(0)}\left(\lVert\phi^{\prime}(x)\lVert_{op}+\lVert\phi(x)\lVert_{\infty}\right),\quad\text{for all }\phi\in C^{r}(B_{5}(0),\mathbb{R}^{n}).$
For $10<L<$ 1 we set the open ball
$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need an $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
[…]
for $[\gamma]\in B_{L}$ and $x\in B_{3}(0)$.
We define the operator $T:B_{L}\times B_{3}(0)\times C([0,1],B_{4}(0))\to C([0,1],B_{4}(0))$ by
[…]
By Lemma 4.12, the map $T$ is $C^{\infty,\infty,r}$ and for each $[\gamma]\in B_{L}$ and $x\in B_{3}(0)$ we have
$\text{Lip}(\Psi([\gamma],x,\cdot))\leq L.$
Thus, by Banach’s Fixed Point Theorem, for each $([\gamma],x)\in B_{L}\times B_{3}(0)$ the map $\Psi([\gamma],x,\cdot)$ has a unique fixed point $\zeta_{[\gamma],x}\in C([0,1],B_{4}(0))$ and
$F:B_{L}\times B_{3}(0)\to C([0,1],B_{4}(0)),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
is $C^{r}$ (see [14, Lemma 6.2]). By the exponential law, we have
  • a)
    $F_{2}:\left(B_{L}\times B_{3}(0)\right)\times[0,1]\to B_{4}(0),\quad([\gamma],t,x)\mapsto\zeta_{[\gamma],x}(t)$ is $C^{r,0}$.
  • b)
    $F_{3}:B_{L}\times[0,1]\to C^{r}(B_{3}(0),B_{4}(0)),\quad([\gamma],t)\mapsto\zeta_{[\gamma]}(t)$ is $C^{r,0}$.
  • c)
    $F_{4}:B_{L}\to C([0,1],C^{r}(B_{3}(0),B_{4}(0))),\quad[\gamma]\mapsto\zeta_{[\gamma]}$ is continuous.
Let $\rho_{2}:C^{r}(B_{3}(0),B_{4}(0))\to C^{r}(B_{2}(0),B_{4}(0))$, $\varphi\mapsto\varphi|_{B_{2}(0)}$, we denote
$H:=C([0,1],\rho_{2})\circ F_{4}:B_{L}\to C([0,1],C^{r}(B_{2}(0),B_{4}(0)),\quad[\gamma]\mapsto\rho_{2}\circ\zeta_{[\gamma]}.$
Then $H$ is continuous. We will show that $H$ has absolutely continuous values.
Let consider the open set
$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{m}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
Then the map
$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{m}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
is continuous. Moreover, since $S$ is linear in the first variable, by Lemma 4.12, the map
$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})$ $\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m}))$
$(\alpha,\beta)$ $\mapsto S\circ(\alpha,\beta)$
is continuous. Hence, for $[\gamma]\in B_{L}$ we have
$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m})).$
This allow us to define the $L^{p}$-absolutely continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{m})$ given by
[…]
If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{m})\to\mathbb{R}^{m}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
$\varepsilon_{x}\left(\tau(t)\right)=\varepsilon_{x}\left(H([\gamma])(t)\right),\quad\text{for all }t\in[0,1].$
Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{m})$, we have $\tau=H([\gamma])$.
Hence $H[\gamma]$ is absolutely continuous and we can consider the map
$H_{AC}:B_{L}\to AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)),\quad[\gamma]\mapsto H([\gamma])$
which is the evolution map for $C^{r}(B_{2}(0),B_{4}(0))$.
Now we will see the case for $L^{p}([0,1],\mathcal{X}^{r}(M))$. We define the 00-neighborhood of $L^{p}([0,1],\mathcal{X}^{r}(M))$
$\mathcal{B}_{L}$ $:=R_{5}^{-1}\left(\prod_{i=1}^{n}B_{L}\right)$
=$=\{[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M)):(\forall i\in\{1,...,n\})\text{ }[\gamma_{i}]\in B_{L}\}$
which is open by continuity of $R_{5}$.
Let $[\gamma]\in\mathcal{B}_{L}$, with $[\gamma_{i}]\in B_{L}$ for each $i\in\{1,...,n\}$. If $m\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $m\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$. Let $\zeta_{[\gamma],\kappa_{j}(m)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
[…]
Then $\zeta_{[\gamma],\kappa_{j}(m)}$ is a $AC_{L^{p}}$-Carathéodory solution to
$x^{\prime}(t)$ $=\gamma_{j}(t)\left(x(t)\right),\quad t\in[0,1]$
$x(0)$ $=\kappa_{j}(m).$
Hence the function
$\eta_{[\gamma],m}:[0,1]\to M,\quad\eta_{[\gamma],m}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(m)}(t)$
is a $AC_{L^{p}}$-Carathéodory solution of the equation
$y^{\prime}(t)$ $=\gamma(t)\left(y(t)\right),\quad t\in[0,1]$
$y(0)$ $=m.$
Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 6.3 the solution of this equation is unique and $\eta_{[\gamma],m}$ is well defined. Moreover, $\tilde{\gamma}$ admits a global flow for initial time $t_{0}=0$, given by
$F_{t,0}^{\overline{\gamma}}(m)=\eta_{[\gamma],m}(t)$
for all $t\in[0,1]$, $m\in M$.
Following point 11.16 of [14], for each $i\in\{1,...,n\}$ we can construct an exponential map
$\text{exp}_{i}:\mathcal{D}_{i}\subseteq TB_{5}(0)\to B_{5}(0)$
such that
$\kappa_{i}^{-1}\circ\text{exp}_{i}=\text{exp}\circ T\kappa_{i}^{-1}|_{\mathcal{D}_{i}}.$
There is an open set $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ containing $\overline{B_{4}(0)}\times\{0\}$ such that $(\text{pr}_{1},\exp_{i})(Q_{i})$ is open in $B_{5}(0)\times B_{5}(0)$ and the map
$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0)$
is a $C^{\infty}$-diffeomorphism onto its image. Assuming that
$T\kappa_{i}^{-1}(\mathcal{O}_{i})\subseteq\mathcal{W},$
since
$\{(x,x):x\in\overline{B_{5}(0)}\}\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i}),$
there exists $s_{i}\in]0,1]$ such that
$\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i})$
is a smooth diffeomorphism and there exists an $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
$(\text{Id}_{B_{4}(0)},\theta_{i}):\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq B_{5}(0)\times B_{5}(0)\to\psi_{i}^{-1}\left(\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\right)\subseteq\mathcal{O}_{i}$
with open image in $\mathcal{O}_{i}$. For each $x\in B_{4}(0)$ we define the set
$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{m}:(x,y)\in\mathcal{O}_{i}\}.$
Hence, we have
$\theta_{i}(x,\cdot)=\left(\exp(x,\cdot)|_{\mathcal{O}_{i,x}}\right)^{-1}|_{B_{s_{i}}(0)}.$
The set
$Z_{i,x}:=\left\{\varphi\in C^{r}(B_{2}(0),\mathbb{R}^{n}):(\forall x\in\overline{B_{1}(0)})\text{ }\varphi(x)\in B_{s_{i}}(x)\right\}$
is open in $C^{r}(B_{2}(0),\mathbb{R}^{m})$ and the map
$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{m})\to C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
is smooth since $\theta$ is smooth [16, Proposition 4.23].
Let $\Psi:\mathcal{U}\to\mathcal{V}$ be a chart on $\text{id}_{M}\in\text{Diff}^{r}(M)$, then since $\rho_{1}$ is a topological embedding we can assume that
$\mathcal{V}:=\rho_{1}^{-1}\left(\prod_{i=1}^{n}\mathcal{V}_{i}\right)$
for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{m})$.
Since $(\theta_{i})_{*}(\text{id}_{B_{2}(0)})=0$, by continuity of $(\theta_{i})_{*}$ there exists open $\text{id}_{B_{2}(0)}$-neighborhoods $\mathcal{Y}_{i}\subseteq Z_{i,x}$ such that
$(\theta_{i})_{*}(\mathcal{Y}_{i})\subseteq\mathcal{V}_{i}.$
Since $H(0)(t)=\text{id}_{B_{2}(0)}$ for all $t\in[0,1]$, by continuity of $H$ there exists open 00-neighborhoods $\mathcal{P}_{i}\subseteq\mathcal{B}_{L}$ such that
$H(\mathcal{P}_{i})\subseteq C([0,1],\mathcal{Y}_{i})$
If
$\mathcal{P}:=R_{5}^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
then $\mathcal{P}$ is an open 00-neighborhood in $L^{p}([0,1],\mathcal{X}^{r}(M))$ with $\mathcal{P}\subseteq\mathcal{B}_{L}$.
This allow us to do the composition
$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})$
Following the same steps of point $1.18$ and $1.19$ of [14], we have that for each $[\gamma]\in\mathcal{P}$, there exists an unique $\theta_{[\gamma]}\in C([0,1],\mathcal{X}^{r}(M))$ such that
$R_{1}(\theta_{[\gamma]})=\Big{(}(\theta_{i})_{*}(H([\gamma_{i}]))\Big{)}_{i=1}^{n}.$
Then $\rho_{1}\left(\theta_{[\gamma]}(t)\right)\in\prod_{i=1}^{n}\mathcal{V}_{i}$, whence $\theta_{[\gamma]}(t)\in\mathcal{V}$. In consequence
$\Psi^{-1}(\theta_{[\gamma]}(t))=\exp\circ\theta_{[\gamma]}(t)\in\text{Diff}^{\text{ }r}(M),\quad\text{ for all }t\in[0,1].$
Let $m\in M$ with $x:=\kappa_{i}(m)$, then we have
$\exp\circ\theta_{[\gamma]}(t)(m)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
$=\kappa_{i}^{-1}\circ\exp_{i}\circ(\theta_{i})_{*}(H([\gamma_{i}]))(t)(x)$
$=\kappa_{i}^{-1}\circ\exp_{i}\circ\left(\exp_{i}|_{\mathcal{W}_{i}}\right)^{-1}\circ H([\gamma_{i}])(t)(x)$
$=\kappa_{i}^{-1}\circ H([\gamma_{i}])(t)(x)$
$=\kappa_{i}^{-1}\circ\zeta_{[\gamma_{i}],x}(t)$
$=F_{t,0}^{\tilde{\gamma}}(m).$
Hence, we define
$\eta_{[\gamma]}:=\exp\circ\theta_{[\gamma]}:[0,1]\to\text{Diff}^{\text{ }r}(M).$
We will show that $\eta_{[\gamma]}$ is absolutely continuous. Since the exponential map $\exp$ is smooth, we need to show that $\theta_{[\gamma]}$ is absolutely continuous, but this is equivalent to show that
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t).$
But this map coincides with the map
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t)$
which is absolutely continuous. Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Proposition 3.8, the composition and hence the product are absolutely continuous. Then we have that
$\theta_{[\gamma]}\in AC_{L^{p}}([0,1],\mathcal{X}^{r}(M))$
whence
$\eta_{[\gamma]}\in AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)).$
In consequence, by Lemma 4.9 and Remark 6.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map
$B_{L}\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad[\gamma_{i}]\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}$
is continuous, the restricted evolution map with continuous values is given by
$\operatorname{Evol}_{C}|_{\mathcal{P}}:\mathcal{P}\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\eta_{[\gamma]}=\exp\circ\theta_{[\gamma]}.$
Hence, by Lemma 4.11 the evolution map $\operatorname{Evol}_{C}$ is continuous and since the restricted right action of $\text{Diff}^{\text{ }r}(M)$ is continuous, by Theorem 4.14, the evolution map
$\operatorname{Evol}:L^{p}([0,1],\mathcal{X}^{r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
Now we will focus in the $L^{p}$-semiregularity of the case $\text{Diff}^{r}(M)$ with $M$ a compact smooth manifold with boundary.
Definition 6.6
Let $M$ and $N$ be smooth manifolds with boundaries and suppose ${f:\partial M\to\partial N}$ is a diffeomorphism, We define adjunction space $M\cup_{f}N$ as the set formed identifying each $p\in\partial M$ with $f(p)\in\partial N$.
We recall [26, Theorem 9.29].
Theorem 6.7
Let $M$ and $N$ be a smooth manifolds with boundaries and ${f:\partial M\to\partial N}$ be a diffeomorphism. Then the adjunction space $M\cup_{f}N$ is a topological manifold without boundary which has a smooth manifold structure such that there are regular domains $M^{\prime},N^{\prime}\subseteq M\cup_{f}N$ diffeomorphic to $M$ and $N$, respectively, such that $M^{\prime}\cup N^{\prime}=M\cup_{f}N$ and ${M^{\prime}\cap N^{\prime}=\partial M^{\prime}=\partial N^{\prime}}$. Moreover, $M$ and $N$ are both compact if and only if $M\cup_{f}N$ is compact.
Definition 6.8
Let $M$ be a smooth manifold with boundary. If $M^{\prime}$ denotes a copy of $M$, we define the double of $M$ as the smooth manifold without boundary
$DM=M\sqcup_{\text{id}_{\partial}}M^{\prime}$ (6.4)
where $\text{id}_{\partial}:\partial M\to\partial M^{\prime}$ is the identity map.
For $p\in M$, we denote by $(p,0)$ and $(p,1)$ the corresponding element in $M$ and $M^{\prime}$, respectively, and if $p\in\partial M$, then $(p,0)\sim(p,1)$. By Theorem 6.7, the map
$\rho:M\to DM,\quad p\mapsto[(p,0)]$
is an embedding onto a regular domain of $DM$ which we identify with $M$.
Definition 6.9
Let $r\in\mathbb{N}\cup\{0,\infty\}$ and $M$ be a smooth manifold with boundary. We define the vector space of $C^{r}$-stratified vector field on $M$ as
$\mathcal{X}_{str}^{r}(M):=\{X\in\mathcal{X}^{r}(M):X(\partial M)\subseteq T\partial M\}.$
We recall [17, Colorally 1.8].
Proposition 6.10
For each $k\in\mathbb{N}\cup\{0,\infty\}$, $n\in\mathbb{N}$, $m\in\{0,...,d\}$ and locally convex space $F$, the restriction map
$\mathcal{E}:C^{k}(\mathbb{R}^{d},F)\to C^{k}([0,\infty)^{m}\times\mathbb{R}^{d-m},F)$
has a continuous linear right inverse. Moreover, the restriction map
$C^{k}(\mathbb{R}^{d},F)\to C^{k}([0,1]^{d},F)$
has continuous linear right inverse.
Proposition 6.11
Let $r\in\mathbb{N}\cup\{0,\infty\}$, $m\in\mathbb{N}$ and $M$ be an $m$-dimensional compact smooth manifold with boundary. Then there exists an extension map
$\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM),\quad X\mapsto\alpha(X)$
which is continuous linear.
Proof.
By compactness of $M$, we find charts $\varphi_{i}:U_{i}\to V_{i}$ of $M$ around points $p_{i}\in\partial M$ such that $(U_{i})_{i=1}^{k}$ is a finite open cover of $\partial M$ which extends to charts $\widetilde{\varphi_{i}}:\widetilde{U_{i}}\to\widetilde{V_{i}}$ around $p_{i}$. For $X\in\mathcal{X}_{str}^{r}(M)$, we write
$Y_{i}:=d\varphi_{i}\circ X\circ\varphi_{i}^{-1}:V_{i}\subseteq[0,\infty)\times\mathbb{R}^{m-1}\to\mathbb{R}^{m}.$
Without loss of generality, we assume that $V_{i}=[0,\infty)\times\mathbb{R}^{m-1}$ and we consider the extension $\widetilde{Y}_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ of $Y_{i}$ given by Proposition 6.10. We define
$\widetilde{X}_{i}:=T\widetilde{\varphi_{i}}^{-1}\circ\widetilde{Y}_{i}\circ\widetilde{\varphi_{i}}:\widetilde{U_{i}}\to T\widetilde{U_{i}}.$
Then $\widetilde{X}_{i}\in\mathcal{X}^{r}(\tilde{U_{i}})$ and the map
$\Phi_{1}:\mathcal{X}_{str}^{r}(M)\to\prod_{i=1}^{k}\mathcal{X}^{r}(\widetilde{U_{i}}),\quad X\mapsto(\widetilde{X}_{i})_{i=1}^{k}$
is continuous linear. Let us consider $\widetilde{U}_{k+1}:=M^{\circ}$, $\widetilde{U}_{k+2}:=(M^{\prime})^{\circ}$ and the open cover of $DM$
$\mathcal{A}:=\{\widetilde{U_{1}},...,\widetilde{U_{k}},\widetilde{U}_{k+1},\widetilde{U}_{k+2}\}.$
Then there exists a partition of the unity subordinate to $\mathcal{A}$, denoted by $\{h_{1},...,h_{k},h_{k+1},h_{k+2}\}$, such that $\text{supp}(h_{i})\subseteq\widetilde{U_{i}}$ for each $i\in\{1,...,k,k+1,k+2\}$.
Denoting the space of all $C^{k}$-vector fields of $\widetilde{U}_{i}$ with support on $\text{supp}(h_{i})$ as $\mathcal{X}_{i}^{r}(\widetilde{U}_{i})$, we see that the map
$\mathcal{X}^{r}(\widetilde{U}_{i})\to\mathcal{X}_{i}^{r}(\widetilde{U}_{i}),\quad Y\mapsto h_{i}Y$
is continuous linear.
Moreover, if $Z\in\mathcal{X}_{i}^{r}(\widetilde{U}_{i})$, we can obtain an extension $\mathcal{E}_{i}(Z)\in\mathcal{X}^{r}(DM)$ of $Z$ by extending it by 00, and a continuous linear map
$\mathcal{E}_{i}:\mathcal{X}_{i}^{r}(\widetilde{U}_{i})\to\mathcal{X}^{r}(DM),\quad Z\mapsto\mathcal{E}_{i}(Z).$
For $X\in\mathcal{X}_{str}^{r}(M)$, we write $\widetilde{X}_{k+1}:=X|_{M^{\circ}}$. This enables us to define the extension map
$\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM),\quad X\mapsto\alpha(X):=\sum_{i=1}^{k+1}\mathcal{E}_{i}(h_{i}\widetilde{X}_{i})$
which is continuous linear. ∎
Remark 6.12
Let $M$ be a compact smooth manifiold with boundary and $r\in\mathbb{N}\cup\{\infty\}$. By [20, Proposition 1.3], the set $\text{Diff}^{r}(M)$ is a embedded submanifold of $C^{r}(M,DM)$. This allows us to consider the inclusion map restricted onto its image
$J:\text{Diff}^{r}(M)\to J(\text{Diff}^{r}(M))\subset C^{r}(M,DM),\quad\phi\mapsto\phi$ (6.5)
then $J$ is a diffeomorphism. Since for each $g\in C^{r}(M,M)$ fixed, the right translation map
$\rho_{C^{r}}(g):C^{r}(M,DM)\to C^{r}(M,DM),\quad\phi\mapsto\phi\circ g$
is smooth. For each $g\in\text{Diff}^{r}(M)$ fixed, the right translation map
$\rho(g):\text{Diff}^{r}(M)\to\text{Diff}^{r}(M),\quad\phi\mapsto\phi\circ g$
can be written as
$\rho(g)=J^{-1}\circ\rho_{C^{r}}(J(g))\circ J.$
Hence $\text{Diff}^{r}(M)$ is a right half-Lie group.
Theorem 6.13
Let $M$ be a compact smooth manifold with boundary, $r\in\mathbb{N}\cup\{\infty\}$ and $1\leq p<\infty$. Then the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular and the evolution map with continuous values
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}\text{Diff}^{r}(M))\to C([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma)$
is continuous.
Proof.
Let consider the map $\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM)$ as Proposition 6.11, we define the map
$\tilde{\alpha}:=L^{p}([0,1],\alpha):L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to L^{p}([0,1],\mathcal{X}^{r}(DM)),\quad[\gamma]\mapsto[\alpha\circ\gamma];$
which is linear and continuous. Since $DM$ is a compact smooth manifold without boundary, the right half-Lie group $\text{Diff}^{r}(DM)$ is $L^{p}$-semiregular with continuous evolution map denoted by
$\operatorname{Evol}_{DM}:L^{p}([0,1],\mathcal{X}^{r}(DM))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(DM)),\quad[\gamma]\mapsto\operatorname{Evol}_{DM}(\gamma).$
Let $[\gamma]\in L^{p}([0,1],\mathcal{X}_{str}^{r}(M))$, we define the absolutely continuous function ${\xi_{\gamma}:[0,1]\to\text{Diff}^{r}(DM)}$ by $\xi_{\gamma}:=\operatorname{Evol}_{DM}(\tilde{\alpha}([\gamma]))$. Then $\xi_{\gamma}$ is the solution of the equation
$\dot{\xi}_{\gamma}(t)$ $=\tilde{\alpha}([\gamma])(t).\xi_{\gamma}(t),\quad t\in[0,1]$
$\xi_{\gamma}(0)$ $=e.$
For $[\gamma]$ close to 00, the proof of Theorem 6.5 shows that, for each $p\in M$, the function $x_{p}:[0,1]\to DM$, given by ${x_{p}(t):=\xi_{\gamma}(t)(p)}$ is a solution of the equation
$\dot{x_{p}}(t)$ $=\tilde{\alpha}([\gamma])(t)\circ x_{p}(t),\quad t\in[0,1]$
$x_{p}(0)$ $=p,$
and this ODE satisfies local uniqueness of Caratheodory solutions. Looking at the compact manifold $\partial M$ without boundary and the vector fields $\gamma(t)|_{\partial M}\in\mathcal{X}^{r}(\partial M)$ we likewise get a solution $y_{p}:[0,1]\to\partial M$ for each $p\in\partial M$, for the differential equation
$\dot{y_{p}}(t)$ $=\tilde{\alpha}([\gamma])(t)\circ y_{p}(t),\quad t\in[0,1]$
$y_{p}(0)$ $=p.$
Then $y_{p}$ also solves the initial value problem for $x_{p}$, whence $x_{p}=y_{p}$ by local uniqueness. In consequence $x_{p}([0,1])\subseteq\partial M$. This implies that for each $t\in[0,1]$ fixed, we have $\xi_{\gamma}(t)(M)\subseteq M$, $\xi_{\gamma}(t)(M^{\prime})\subseteq M^{\prime}$ and $\xi_{\gamma}(t)(\partial M)\subseteq\partial M$. Therefore, we obtain ${\xi_{\gamma}(t)|_{M}\in\text{Diff}^{r}(M)}$.
Consider the smooth embedding $\iota:\text{Diff}^{r}(DM)\to C^{r}(DM,DM)$, $\phi\mapsto\phi$. Then the map
$\tilde{\iota}:=AC_{L^{p}}([0,1],\iota):AC_{L^{p}}([0,1],\text{Diff}^{r}(DM))\to AC_{L^{p}}([0,1],C^{r}(DM,DM)),\quad\eta\mapsto\iota\circ\eta$
is smooth. Let $\rho:M\to DM$ be the inclusion map, which is smooth and a diffeomorphism onto its image, then the maps
$\rho^{*}:=C^{r}(\rho,DM):C^{r}(DM,DM)\to C^{r}(M,DM),\quad\phi\mapsto\phi\circ\rho=\phi|_{M}$
and
$\tilde{\rho}:AC_{L^{p}}([0,1],C^{r}(DM,DM))\to AC_{L^{p}}([0,1],C^{r}(M,DM)),\quad\eta\mapsto\rho^{*}\circ\eta$
are smooth.
Consider the restricted inclusion map $J$ as in equation (6.5), by Lemma 2.25, we define the map
$\tilde{j}:AC_{L^{p}}([0,1],J(\text{Diff}^{r}(M)))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(M)),\quad\eta\mapsto J^{-1}\circ\eta.$
The fact that $\xi_{\gamma}(t)|_{M}\in\text{Diff}^{r}(M)$ for each $t\in[0,1]$, enables us to define the function
$\zeta_{\gamma}:[0,1]\to C^{r}(M,DM),\quad\zeta_{\gamma}(t):=(\tilde{j}\circ\tilde{\rho}\circ\tilde{\iota})(\xi_{\gamma})(t)$
which is in $AC_{L^{p}}([0,1],\text{Diff}^{r}(M))$. Moreover, by definition of $\alpha$, we have
$\tilde{\alpha}([\gamma])(t)\circ\zeta_{\gamma}(t)=\gamma(t)\circ\zeta_{\gamma}(t).$
Hence $\zeta_{\gamma}$ verifies the equation
$\dot{\zeta_{\gamma}}(t)$ $=\gamma(t).\zeta_{\gamma}(t),\quad t\in[0,1]$
$\zeta_{\gamma}(0)$ $=e.$
Therefore $\text{Diff}^{r}(M)$ is $L^{p}$-semiregular and the evolution map is given by
$\operatorname{Evol}:L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\Big{(}\tilde{j}\circ\tilde{\rho}\circ\tilde{\iota}\circ\operatorname{Evol}_{DM}\circ\tilde{\alpha}\Big{)}([\gamma]).$
We consider the inclusion map
$\omega:AC_{L^{p}}([0,1],C^{r}(M,DM))\to C([0,1],C^{r}(M,DM)),\quad\eta\mapsto\eta$
which is smooth, and we define the smooth map
$\widetilde{\rho}_{C}:AC_{L^{p}}([0,1],C^{r}(DM,DM))\to C([0,1],C^{r}(M,DM)),\quad\eta\mapsto\omega(\rho^{*}\circ\eta).$
We write
$\widetilde{j}_{C}:C([0,1],J(\text{Diff}^{r}(M)))\to C([0,1],\text{Diff}^{r}(M)),\quad\eta\to J^{-1}\circ\eta$
then $\widetilde{j}_{C}$ is continuous. Therefore, the evolution map with continuous values $\operatorname{Evol}_{C}$ is given by
$\operatorname{Evol}_{C}:L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to C([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\Big{(}\widetilde{j}_{C}\circ\widetilde{\rho}_{C}\circ\widetilde{\iota}\circ\operatorname{Evol}_{DM}\circ\widetilde{\alpha}\Big{)}([\gamma])$
and is continuous since is composition of continuous maps. ∎
Acknowledgements: The author would like to thank Helge Glöckner for his guidance during the development of this work. The author was partially supported by the FONDECYT Grant #1241719 and by ANID and DAAD (DAAD/BecasChile 2020, ID:91762237/62190017).